### v1

# Enhancing Fault Localization Through Ordered Code Analysis with LLM Agents and Self-Reflection

Abstract.
Locating and fixing software faults is a time-consuming and resource-intensive task in software development. Traditional fault localization methods, such as Spectrum-Based Fault Localization (SBFL), rely on statistical analysis of test coverage data but often suffer from lower accuracy. Learning-based techniques, while more effective, require extensive training data and can be computationally expensive. Recent advancements in Large Language Models (LLMs) offer promising improvements in fault localization by enhancing code comprehension and reasoning. However, these LLM-based techniques still face challenges, including token limitations, degraded performance with long inputs, and difficulties managing large-scale projects with complex systems involving multiple interacting components. To address these issues, we introduce LLM4FL, a novel LLM-agent-based fault localization approach that integrates SBFL rankings with a divide-and-conquer strategy. By dividing large coverage data into manageable groups and employing multiple LLM agents through prompt chaining, LLM4FL navigates the codebase and localizes faults more effectively. The approach also incorporates self-reflection and chain-of-thought reasoning, enabling agents to iteratively generate fixes and re-rank suspicious methods. We evaluated LLM4FL on the Defects4J (V2.0.0) benchmark, comprising 675 real-world faults from 14 open-source Java projects. Our results demonstrate that LLM4FL outperforms AutoFL by 19.27% in Top-1 accuracy and surpasses state-of-the-art supervised techniques such as DeepFL and Grace, all without task-specific training. Additionally, we highlight the impact of coverage splitting and prompt chaining on fault localization performance and show that different method ordering can improve Top-1 accuracy by up to 22%.
1. Introduction
The process of locating and fixing software faults requires significant time and effort. Research shows that software development teams allocate more than half of their budgets to testing and debugging activities (Hait and Tassey, 2002; Alaboudi and LaToza, 2021). As software systems become increasingly complex, the demand for more efficient and accurate fault localization techniques continues to grow. Hence, to assist developers and reduce debugging costs, researchers have developed various fault localization techniques (Li et al., 2019; Lou et al., 2021; Qian et al., 2023b; Li et al., 2021; Abreu et al., 2009; Sohn and Yoo, 2017). These techniques analyze code coverage and program execution to identify the most likely faulty components, assisting developers in finding the fault.
However, despite the advances in fault localization techniques, many existing approaches still struggle with scalability and precision. Traditional methods, such as Spectrum-Based Fault Localization (SBFL), use statistical analysis to analyze coverage data from passing and failing test cases to rank suspicious code elements (Abreu et al., 2006). While these techniques provide valuable insights, their accuracy is lower. Their reliance on statistical correlations between test failures and code coverage does not always capture the deeper semantic relationships needed for more accurate fault localization (Wong et al., 2016; Xie et al., 2013; Le et al., 2013). To address these issues, recent techniques applied machine learning and deep learning models to improve fault localization (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021, 2019). These methods enhance the ranking of suspicious code elements by incorporating additional information like code complexity, text similarity, and historical fault data. Researchers also leveraged models like Graph Neural Networks (GNNs) to represent code structures and achieved state-of-the-art fault localization accuracy (Lou et al., 2021; Qian et al., 2023b; Rafi et al., 2024; Lou et al., 2021). However, these techniques often require extensive training data and significant training time.
Recent advances in Large Language Models (LLMs) have shown great potential for fault localization due to their strong language comprehension and generation capabilities (Abedu et al., 2024; Lin et al., 2024). LLMs trained on extensive programming datasets can understand code structure, interpret error messages, and even suggest fixes for common software bugs (Kang et al., 2024; Wu et al., 2023; Pu et al., 2023; Li et al., 2023). These models, with their ability to analyze and process both natural language and code, present an opportunity to significantly improve traditional fault localization methods by incorporating deeper semantic analysis and context-aware reasoning. Wu et al. (Wu et al., 2023), directly present LLMs with faulty methods or classes with test failure information and ask to locate the issue, which provides valuable insights. Kang et al. (Kang et al., 2024) which operates as an automated fault localization technique that leverages LLMs to localize fault given a single failing test. It focuses on method-level fault localization and provides bug location and also a natural language explanation of why a particular code location is likely to be faulty.
Despite their potential, existing LLM-based fault localization techniques face several challenges. The token limitations of LLMs restrict their ability to effectively process long code files or large sets of code coverage data, which is often required when dealing with large-scale software systems (Hadi et al., 2023; Hou et al., 2023; Wu et al., 2023). Additionally, the performance of LLMs can degrade when applied to complex systems that require the model to reason over multiple interacting components, making it difficult to maintain accuracy and consistency across broader projects (Levy et al., 2024; Liu et al., 2024). Furthermore, current LLM-based techniques have yet to fully explore how these models can be effectively integrated with traditional fault localization techniques to maximize their strengths in a complementary and efficient manner (Wu et al., 2023; Kang et al., 2024; Yang et al., 2024; Bin Murtaza et al., 2024).
In this paper, we propose LLM4FL, an LLM-based fault localization. To address the challenges of analyzing large-scale software projects, where code coverage and complexity often exceed LLM token limits (Wu et al., 2023; Liu et al., 2024), LLM4FL implements a divide-and-conquer strategy. Before applying the divide-and-conquer strategy, we use an SBFL technique to sort the covered methods, building on findings from prior research (Chen et al., 2024) that demonstrate LLM performance improves when the order of instructions is carefully considered. Then, we divide the coverage data into manageable groups that the LLM can process within its token limits. Note that, regardless of the order, LLM4FL eventually analyzes every covered method.
In addition to using a divide-and-conquer strategy, LLM4FL takes inspiration from how human developers debug software. Developers typically analyze multiple types of information, including error messages, stack traces, and code snippets, to incrementally narrow down potential faulty components (Alaboudi and LaToza, 2021; Böhme et al., 2017). LLM4FL emulates this process by utilizing two LLM agents collaborating to iteratively and autonomously navigate the code to locate the faults. LLM4FL implements a Tester and a Debugger Agent, each tasked with specialized tools to assist in the fault localization process. The Tester Agent identifies and prioritizes suspicious methods by analyzing the failing test, and stack traces in different groups of covered methods to list a list of highly suspicious methods. It mimics the developer’s process of investigating suspicious areas in the code by understanding the context of error messages. Meanwhile, the Debugger Agent thoroughly evaluates the given list of candidate methods by navigating the code and ranks them based on its analysis. These two agents communicate through a prompt chaining mechanism that allows them to share insights and build upon each other’s findings.
We evaluated LLM4FL using the Defects4J (V2.0.0) benchmark (Just et al., 2014), which contains 675 real-world faults from 14 open-source Java projects. Our results demonstrate that LLM4FL surpasses LLM-based technique AutoFL (Kang et al., 2024), by achieving 19.27% higher Top-1 accuracy. Additionally, LLM4FL outperforms supervised techniques such as DeepFL (Li et al., 2019) and Grace (Lou et al., 2021), even without task-specific training. We also analyzed the impact of individual components within LLM4FL on fault localization accuracy. Our findings indicate that each component plays a significant role in its performance, with coverage splitting and prompt chaining contributing the most. When these components are removed, accuracy drops considerably, underscoring their importance in managing token limitations and facilitating efficient multi-agent collaboration. Moreover, we examined whether the initial ordering of methods provided to the LLM influences performance. The results reveal that method ordering is important, with a Top-1 accuracy difference of up to 22% when comparing an execution-based ordering and the order provided by DepGraph (Rafi et al., 2024).
The paper makes the following contributions:
  • We introduce LLM4FL, a novel LLM-based fault localization technique that employs a divide-and-conquer strategy. This technique groups large coverage data and ranks the covered methods using an SBFL formula. By utilizing multiple agents and prompt chaining, LLM4FL navigates the code iteratively to effectively identify and localize faults.
  • LLM4FL demonstrates superior performance, surpassing AutoFL (Kang et al., 2024) by 19.27% in Top-1 accuracy. It also outperforms state-of-the-art supervised techniques like DeepFL and Grace, achieving these results without requiring task-specific training.
  • Our analysis of LLM4FL’s components shows that key features like coverage splitting and prompt chaining are essential to its fault localization accuracy. Removing these features leads to significant performance declines, emphasizing their importance in handling token limitations and enabling effective agent collaboration.
  • We further investigate the effect of method ordering on LLM performance. The study reveals that different ordering can enhance fault localization accuracy by up to 22% in Top-1 scores. While LLM4FLDepGraph achieves the highest overall accuracy, LLM4FLOchiai offers a more efficient solution, balancing accuracy gains with lower computational overhead, making it practical for broader use.
In short, we provide a strategy to mitigate the token limitation issues in LLM-based fault localization and highlight the impact of initial method ordering. The findings may help inspire future research to improve LLM-based fault localization for large-scale software projects.
Paper Organization. Section 2 discusses related work. Section 3 describes our technique, LLM4FL. Section 4 presents the experimental results. Section 5 discusses the threats to validity. Section 6 concludes the paper.
2. Background and Related Work
2.1. Background
Large Language Models. Large Language Models (LLMs), primarily built on the transformer architecture (Meta AI, 2024; Brown, 2020; Roziere et al., 2023), have significantly advanced the field of natural language processing (NLP). These LLMs, such as the widely recognized GPT3 model with its 175 billion parameters (Brown, 2020), are trained on diverse text data from various sources, including source code. The training involves self-supervised learning objectives that enable these models to develop a deep understanding of language and generate contextually relevant and semantically coherent text. LLMs have shown substantial capability in tasks that involve complex language comprehension and generation (Abedu et al., 2024; Lin et al., 2024), such as code recognition and generation. Recent research has leveraged LLMs in software engineering tasks, particularly in fault localization (Kang et al., 2024; Qin et al., 2024; Yang et al., 2024), where they assist in identifying the faulty code groups responsible for software errors. One of the key advantages of using LLMs in fault localization is their ability to process both natural language and code without re-training, allowing them to analyze error messages, stack traces, and test case information to infer suspicious methods or code sections in an unsupervised zero-shot setting.
LLM Agents. LLM agents leverage LLMs to autonomously execute tasks described in natural language, making them versatile tools across various domains. LLM agents are artificial intelligence systems that utilize LLMs as their core computational engines to understand questions and generate human-like responses. They leverage functionalities like memory management (Zhou et al., 2023) and tool integration (Xia et al., 2024; Roy et al., 2024) to handle multi-step and complex operations seamlessly. The agents can refine their responses based on feedback, learn from new information, and even interact with other AI agents to collaboratively solve complex tasks (Hong et al., 2024; Qian et al., 2023a; Xu et al., 2023; Lin et al., 2024). Through prompting, agents can be assigned different roles (e.g., a developer or a tester), providing more domain-specific responses that help improve the answer (Hong et al., 2024; White et al., 2024; Shao et al., 2023). As their potential expands, LLM agents play a crucial role in advancing automation and boosting productivity in software development. In this paper, we explore using LLM agents to improve fault localization by emulating developers’ debugging process.
2.2. Related Work
Spectrum-based Fault Localization. Spectrum-based fault localization (SBFL) (Abreu et al., 2006; Jones et al., 2002; Wong et al., 2013; Abreu et al., 2009) employs statistical techniques to evaluate the suspiciousness of individual code elements, such as methods, by analyzing test outcomes and execution traces. The core idea of SBFL is that code components that are executed more frequently in failing tests and less frequently in passing tests are more likely to contain faults. Despite its widespread study, SBFL’s practical effectiveness remains limited (Kochhar et al., 2016; Xie et al., 2016). To enhance SBFL’s accuracy, recent research (Cui et al., 2020; Wen et al., 2019; Chen et al., 2022; Xu et al., 2020b) has suggested incorporating additional data, such as code changes (Wen et al., 2019; Chen et al., 2022) or mutation analysis (Cui et al., 2020; Xu et al., 2020b). However, SBFL’s reliance on code coverage metrics still poses challenges, as its suspiciousness scores may not generalize effectively to different faults or systems.
Learning-based fault localization. Recent efforts have focused on improving SBFL with learning-based methods (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021; Li and Zhang, 2017; Li et al., 2019; Zhang et al., 2019a). These approaches use machine learning models like radial basis function networks (Wong et al., 2011), back-propagation networks (Wong and Qi, 2009), and convolutional neural networks (Zhang et al., 2019a; Li et al., 2021; Albawi et al., 2017) to estimate suspiciousness scores based on historical faults. Some techniques, such as FLUCCS (Sohn and Yoo, 2017), combine SBFL scores with metrics like code complexity, while others, like DeepFL (Li et al., 2019) and CombineFL (Zou et al., 2019), merge multiple sources such as spectrum-based and mutation-based data (Moon et al., 2014; Papadakis and Le Traon, 2015; Dutta and Godboley, 2021). Graph neural networks (GNNs) have also been applied to fault localization (Qian et al., 2023b; Lou et al., 2021; Qian et al., 2021; Xu et al., 2020a). Techniques like Grace (Lou et al., 2021) and GNET4FL (Qian et al., 2023b) utilize test coverage and source code structure for improved accuracy, while DepGraph (Rafi et al., 2024) refines these approaches by graph pruning and incorporating code change information, resulting in higher performance with reduced computational demands. Although these learning-based techniques show improved results, they require training data that may not be available to every project, and the training process can be expensive due to model complexity.
LLM-Based Fault Localization. Large Language Models (LLMs), such as GPT-4o (OpenAI, 2024), LLaMA (Meta AI, 2024), and ChatGPT (Achiam et al., 2023), demonstrated remarkable abilities in processing both natural and programming languages. LLMs have shown potential in identifying and fixing errors using program code and error logs (Achiam et al., 2023). However, one of the major challenges that LLMs face in fault localization is the token limitation issue. LLMs are restricted to a fixed number of tokens, ranging from 2,000 to 128,000 (OpenAI, 2024; Meta AI, 2024), which poses difficulties in handling large-scale software projects with long stack traces and extensive codebases. This limitation can lead to incomplete analyses, as critical context might be truncated or lost, forcing models to work with fragmented information, ultimately affecting the overall fault localization performance (Hadi et al., 2023; Hou et al., 2023).
Moreover, LLM-based fault localization techniques often focus on localizing faults within small code snippets due to these context limitations. For example, LLMAO (Yang et al., 2024) uses lightweight bidirectional adapters on LLMs to generate suspiciousness scores for code lines, but only within a limited context of 128 lines of code. Similarly, Wu et al. (Wu et al., 2023) prompt ChatGPT with code and error logs to identify faulty lines, but these methods struggle to scale to larger software projects (Liu et al., 2024). AutoFL (Kang et al., 2024) is an LLM-based fault localization technique that detects faulty code locations and generates natural language explanations for bugs. However, it also faces difficulty handling large-scale projects due to LLM context length constraints and reduced performance on complex bugs requiring deeper repository exploration. To address these challenges, there is a growing need for approaches that enable LLMs to localize faults across entire software projects, ensuring they can handle larger inputs while maintaining accuracy and effectiveness. Thus, we propose LLM4FL, that leverages a divide-and-conquer approach to navigate through codebases and identify faults through multiple-agent collaboration.
3. Methodology
Figure 1. An overview of LLM4FL.
Applying Large Language Models (LLMs) for fault localization presents specific challenges, such as token size limitations and performance issues with longer input contexts (Levy et al., 2024; Liu et al., 2024; Wu et al., 2023). Processing large datasets like test coverage and source code in a single query is impractical due to these constraints. Moreover, there are limited studies on effective prompting techniques for fault localization. To address these issues, we propose LLM4FL, an LLM-agent-based fault localization system. Figure 1 illustrates an overview of our technique. LLM4FL operates in four stages: 1) Analyzing the Covered Methods Using a Divide-and-Conquer Strategy, 2) Defining LLM Agent Roles and Tools, 3) Code Navigation through Agent Collaboration and Prompt Chaining, and 4) Self-Reflection and Chain-of-Thought Reasoning. First, the coverage data is divided into smaller, manageable groups to fit within the LLM’s token limits. Second, LLM agents are assigned specific roles and equipped with tools to analyze these groups. Third, LLM agents collaborate through prompt chaining to navigate the code autonomously and identify potential faults. Finally, the agents refine their analysis using self-reflection and chain-of-thought reasoning. Below, we provide more details on each of these stages.
3.1. Analyzing the Covered Methods Using Divide-and-Conquer
Prior studies in fault localization utilized dynamic execution data, such as coverage information from both failing and passing test cases, to identify faulty code locations (Lou et al., 2021; Li et al., 2019; Abreu et al., 2009; Rafi et al., 2024; Qian et al., 2023b). However, Large Language Models (LLMs) face challenges when processing large-scale coverage data due to inherent token size limitations (Kang et al., 2024). These limitations lead to performance degradation, extended processing times, and reduced accuracy when input tokens are truncated. To address these constraints, we propose a divide-and-conquer strategy, dividing the coverage data (i.e., containing the covered methods and the corresponding code) into manageable groups that the LLM can process within its token limits.
We generate a list of covered methods using tools like GZoltar (Campos et al., 2012), focusing only on methods executed by failing tests. To manage the token limitation issue, we include only the specific method statements that were executed by the failing tests, excluding any uncovered portions of the code. Before doing a divide-and-conquer, we use a Spectrum-Based Fault Localization (SBFL) technique to sort the covered methods so that those more likely to be faulty can be grouped together. This approach is inspired by a recent study (Tyen et al., 2024) that demonstrates LLMs perform better when the order of instruction is carefully considered. Note that, regardless of the order, LLM4FL eventually analyzes every covered method. Specifically, we use Ochiai (Abreu et al., 2006) to sort the methods before the divide-and-conquer process. Ochiai is a widely used unsupervised spectrum-based fault localization technique because of its high efficiency and good fault localization accuracy (Abreu et al., 2006; Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021). Intuitively, Ochiai assigns a higher suspiciousness score to a statement that is executed more frequently by failing cases and less frequently by passing test cases. Ochiai is defined as:
$Ochiai(a_{ef},a_{nf},a_{ep})=\frac{a_{ef}}{\sqrt{(a_{ef}+a_{nf})\times(a_{ef}+a_{ep})}},$
where $a_{ef}$ is the number of failed test cases that execute a statement, $a_{nf}$ is the number of failed test cases that do not execute the statement, and $a_{ep}$ is the number of passed test cases that execute the statement. The result of Ochiai is a value between 0 and 1, with higher values indicating a higher likelihood that a particular statement is faulty. We then aggregate the scores for every statement in a method to get the method-level score.
Once the covered methods are sorted, we divide the methods into smaller groups, ensuring that each group fits within the token limitation of the LLM. More formally, let $C$ represent the sorted covered methods, containing a sequence of pairs ($m$, $s$), where $m$ denotes a covered source method and $s$ denotes the corresponding code statements. The group length is determined using the token limitation specified in the official documentation of GPT-4o-mini (the LLM that we use in this study), which has an input context window of 128,000 tokens (OpenAI, 2024). The length is calculated as $TotalTokenLength/TokenLimitation$, ensuring each group remains within the model’s token limitation. Specifically, we divide $C$ into to sequence of $C_{1},C_{2},\ldots,C_{n}$ each $C_{i}$ contains the subset of pairs ($m_{i}$, $s_{i}$), and satisfies the constraint —$C_{i}$$<=TokenLimitation$.
3.2. LLM Agents and the Available Tools
LLM4FL emulates developers’ debugging process by implementing two LLM agents: 1) Tester and 2) Debugger. Below we discuss the details of these agents:
Tester Agent. The tester agent’s main goal is to identify and prioritize suspicious methods. More formally, it navigates through each group $C_{i}$ of the covered methods. The Tester Agent’s input includes the sorted covered methods in each group $C_{i}$ and the corresponding test code, stack traces, and related test information from failing executions. To further reduce the input token size and allow LLM to focus on the relevant code, we only provide the portion of the test code leading to the failure as input. We analyze the stack trace to identify where the test failed and extract the code up to that point, omitting the rest.
Figure 2 gives an example of the Tester Agent’s prompt. The Tester Agent is responsible for identifying suspicious methods likely related to faults in the system. To achieve this, it needs to 1) analyze methods executed by failing tests, 2) review relevant test code, and 3) examine stack trace data to pinpoint where failures occurred. To facilitate this process, we provide the Tester Agent with three tools. The tool get_covered_methods allows the agent to retrieve the source code of the methods executed during failing tests, get_test_code fetches the relevant test code, and get_stacktrace gathers stack trace data. Each tool has a unique name and description, which the agent uses to determine when to apply the tool during its analysis. The agent decides autonomously, based on its task, which tool to invoke by interpreting the tool’s description and relevance to the current stage of fault localization. Additionally, each method within the grouped data is assigned a unique ID during the dividing process, allowing the agent to reference and retrieve specific methods. After analyzing all groups, the Tester Agent generates a list of suspicious methods, denoted as $S_{i}\subseteq M_{i}$, where $M_{i}$ represents the methods in group $C_{i}$. This list of suspicious methods is then consolidated into a final list of method IDs for further analysis.
Prompt Template for Initial Candidate Selection with Advanced Tools As an Intelligent Tester Agent, utilize the following tools to identify methods most likely related to the fault. Focus on analyzing the test code, stack trace, and covered methods to pinpoint suspicious or fault-inducing methods. Ensure the exact method_id is included in the final JSON output. Available Tools:
get_test_code()       – Retrieves the complete test code.
get_stacktrace()         – Access the stack trace of the failed test.
get_covered_methods()   – Lists all methods executed during the test.
Your Task: Use the insights from the tools and provide the potential suspicious methods in the following JSON format:
        "method_ids": [1, 2, 3, ...]  // The potential suspicious method’s ids
Figure 2. Enhanced Prompt for Initial Candidate Selection.
Debugger Agent. Figure 3 gives an example of the Debugger Agent’s prompt. The Debugger Agent is responsible for thoroughly evaluating and ranking a given list of candidate methods $S$. To perform its analysis, the Debugger Agent relies on three tools designed to extract relevant information. get_test_code and get_stacktrace provide LLM’s requested information by retrieving the relevant test code and stack trace data, helping the Debugger Agent understand the failure scenarios associated with each method. Additionally, the get_method_body tool allows the agent to fetch the full source code of a method using its unique method ID, enabling the Debugger Agent to analyze each method’s source code.
Once the Debugger Agent completes its analysis, it assigns a rank score $\sigma_{j}$ to each method $m_{j}$, representing the likelihood that the method is contributing to the fault. The candidate methods in $S$ are then ranked based on these scores, producing a ranked list $R$ that orders the methods from most to least suspicious:
$R=\text{sort}(\{(m_{j},\sigma_{j}):m_{j}\in S\})$
This final ranked list $R$ is output in JSON format for further evaluation and parsing.
Prompt Template for Code Understanding and Fault Analysis As an Advanced Debugger Agent, use the following tools to analyze the provided method bodies based on the Tester Agent’s insights from the test and stack traces. The fault may propagate across multiple methods or be indirectly related to the original issue. Available Tools:
get_test_code()       – Retrieve the entire test code for detailed analysis.
get_stacktrace()         – Access the stack trace to trace method calls and identify potential faults.
get_method_body()      – Retrieve the body of specific methods for further inspection.
Task: Using the provided tools, analyze and focus on the following Methods: {suspicious methods} Rank these methods from most to least suspicious, and provide brief reasoning for each based on its behavior and potential involvement in the fault. IMPORTANT: Provide your output in the following JSON format:
            "method_id": int,  // The most suspicious method’s id
            "reasoning": string,  // Reasoning for its suspiciousness
            "rank": int  // Rank of suspiciousness
Figure 3. Enhanced Prompt for Code Understanding and Fault Analysis.
3.3. Code Navigation Through Prompt Chaining
Code navigation is a critical part of how developers trace faults in real-world scenarios. When debugging, developers often start by analyzing a specific method and then explore its caller or callee methods to better understand the overall logic and pinpoint where the fault might lie. Our approach emulates this process by implementing code navigation through prompt chaining, allowing the Tester and Debugger Agents to collaborate and progressively focus their analysis on the most relevant parts of the codebase.
The process begins with the Tester Agent, which works through each group of the coverage data $C=\{C_{1},C_{2},\ldots,C_{n}\}$, where each group $C_{i}$ contains a subset of the covered methods $M_{i}$. The Tester Agent identifies suspicious methods within each group, producing a list $S_{i}\subseteq M_{i}$ of candidate methods. This step mimics how a developer might flag areas of interest in the code that seem related to the fault. After analyzing all the groups, the Tester Agent compiles the suspicious methods into a consolidated list:
$S=\bigcup_{i=1}^{n}S_{i}$
This list $S$ is passed to the Debugger Agent, initiating the prompt chaining process.
The Debugger Agent takes the output from the Tester Agent and dives deeper into the suspicious methods. In this stage, the Debugger Agent navigates through each method in $S$, inspecting not only the method itself but also the related methods, such as caller and callee methods, if the agent believes further analysis can help identify faulty code location. The Debugger Agent emulates the manual process of tracing how faults propagate through different parts of the code. By leveraging its tools, the Debugger Agent retrieves information on how each method interacts with others and how it may contribute to the fault. If the Debugger Agent identifies that a method $m_{j+1}$ is called within $m_{j}$ and might require further examination, it can request the code for $m_{j+1}$ if it is part of the candidate list, ensuring deeper analysis where necessary. LLM4FL implements an LLM-based fault navigation system that allows the Debugger Agent to dynamically traverse the codebase and retrieve additional relevant methods for further inspection. Through this process, the Debugger Agent traces potential fault propagation pathways and evaluates how faults spread across methods.
3.4. Self-Reflection and Chain-of-Thought for Improved Fault Localization Results
LLM4FL uses two additional prompting techniques to further improve the fault localization results.
Self-Reflection. In our approach, both the Tester and Debugger Agents engage in self-reflection to enhance the quality of their outputs. After completing their initial analysis, each agent operates in two phases: a generator phase and a reflector phase. During the generator phase, the agents produce their primary results, such as identifying suspicious methods or ranking them by suspiciousness. In the reflector phase, the agents critique their own work, offering feedback to refine their results and correct potential errors. While it has not yet been studied in the fault localization context, a recent study (Renze and Guven, 2024) found that self-reflection improves LLM’s performance in problem-solving tasks.
Chain-of-Thought for Refining Fault Localization. We incorporate a chain-of-thought reasoning process to enhance fault localization. After generating a ranked list of suspicious methods, the LLMs propose potential fixes for the top-ranked methods. The agents then re-rank the methods by considering the generated fixes. This process allows the agents to think more deeply about the faults and reassess the ranking based on the insights gained from considering these fixes (Bao et al., 2024; Wang et al., 2024).
4. STUDY DESIGN AND RESULTS
In this section, we first describe the study design and setup. Then, we present the motivation, approach, and results of the research questions.
Table 1. An overview of our studied projects from Defects4J v2.0.0. #Faults, LOC, and #Tests show the number of faults, lines of code, and tests in each system. Fault-triggering Tests shows the number of failing tests that trigger the fault.
         Project          #Faults          LOC          #Tests          Fault-triggering Tests
         Cli          39          4K          94          66
         Closure          174          90K          7,911          545
         Codec          18          7K          206          43
         Collections          4          65K          1,286          4
         Compress          47          9K          73          72
         Csv          16          2K          54          24
         Gson          18          14K          720          34
         JacksonCore          26          22K          206          53
         JacksonXml          6          9K          138          12
         Jsoup          93          8K          139          144
         Lang          64          22K          2,291          121
         Math          106          85K          4,378          176
         Mockito          38          11K          1,379          118
         Time          26          28K          4,041          74
         Total          675          380K          24,302          1,486
Benchmark Dataset. To answer the RQs, we conducted the experiment on 675 faults across 14 projects from the Defects4J benchmark (V2.0.0) (Just et al., 2014). Defects4J provides a controlled environment to reproduce faults collected from projects of various types and sizes. Defects4J is widely used in prior automated fault localization research (Lou et al., 2021; Sohn and Yoo, 2017; Chen et al., 2022; Zhang et al., 2017). We excluded three projects, JacksonDatabind, JxPath, and Chart, from Defects4J in our study because we encountered many execution errors and were not able to collect test coverage information for them. Table 1 gives detailed information on the projects and faults we use in our study. In total, the faults have over 1.4K fault-triggering tests (i.e., failing tests that cover the fault). The sizes of the studied projects range from 2K to 90K lines of code. Note that since a fault may have multiple fault-triggering tests, there are more fault-triggering tests than faults.
Evaluation Metrics. According to prior findings, debugging faults at the class level lacks precision for effective location (Kochhar et al., 2016). Alternatively, pinpointing them at the statement level might be overly detailed, omitting important context (Parnin and Orso, 2011). Hence, in keeping with prior work (Benton et al., 2020; B. Le et al., 2016; Li et al., 2019; Lou et al., 2021; Vancsics et al., 2021), we perform our fault localization process at the method level. We apply the following commonly-used metrics for evaluation:
Recall at Top-N. The Top-N metric measures the number of faults with at least one faulty program element (in this paper, methods) ranked in the top N. The result from LLM4FL is a ranked list based on the suspiciousness score. Prior research (Parnin and Orso, 2011) indicates that developers typically only scrutinize a limited number of top-ranked faulty elements. Therefore, our study focuses on Top-N, where N is set to 1, 3, 5, and 10.
Implementation and Environment. To collect test coverage data and compute results for baseline techniques, we utilized Gzoltar (Campos et al., 2012), an automated tool that executes tests and gathers coverage information. For the LLM-based components, we employed OpenAI’s GPT-4o mini, a more cost-effective and capable alternative to GPT-3.5 Turbo (OpenAI, 2024). LangChain v0.2 was used to streamline the development of LLM4FL, facilitating the integration of language models with external tools and enhancing the system’s overall functionality (Langchain, 2024a). To implement the self-reflection technique, we leveraged the LangGraph framework, which enabled graph-based reasoning and decision-making processes (Langchain, 2024b). To minimize the variations in the output, we set the temperature parameter to 0 during model inference.
4.1. RQ1: How does LLM4FL perform compared with other fault localization techniques?
Motivation. In this RQ we evaluate the fault localization accuracy of LLM4FL by comparing with various baseline techniques.
Approach. We compare LLM4FL’s fault localization accuracy with five baselines: Ochiai (Abreu et al., 2006), DeepFL (Li et al., 2019), Grace (Lou et al., 2021), DepGraph (Rafi et al., 2024), and AutoFL (Kang et al., 2024).
Ochiai (Abreu et al., 2006) is widely recognized in fault localization research for its high efficiency and accuracy, making it a common baseline for comparison (Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021; Rafi et al., 2024). As such, we use Ochiai to rank the methods during the segmentation process and include it as a baseline for accuracy comparison. DeepFL (Li et al., 2019) is a deep-learning-based fault localization technique that integrates spectrum-based and other metrics such as code complexity, and textual similarity features to locate faults. It utilizes a Multi-layer Perceptron (MLP) model to analyze these varied feature dimensions. We follow the study (Li et al., 2019) to implement DeepFL and include the SBFL scores from 34 techniques, code complexity, and textual similarities as part of the features for the deep learning model. Grace (Lou et al., 2021) is one of the first fault localization techniques based on graph neural networks (GNN) that represents code as a graph and uses a gated graph neural network to rank the faulty methods. Since then, GNN-based techniques have shown superior fault localization accuracy compared to traditional techniques. DepGraph (Rafi et al., 2024) is a GNN-based technique that further improves Grace by enhancing code representation in a graph using interprocedural call graph analysis for graph pruning and integrating historical code change information in the graph.
AutoFL is a LLM-based fault localization technique. It begins by providing the LLM with information about a failing test and descriptions of specific methods that can be used to navigate the codebase. The LLM then interacts with these methods to gather relevant information, such as covered classes, methods, and code snippets. Scoring and ranking candidate methods (depicted as black rectangles) based on five AutoFL prediction outcomes. The methods are ranked by assigning scores based on multiple runs of AutoFL. In each run, a method’s score is the inverse of the total number of predicted methods, and these scores are averaged across all runs. Methods are then ranked in descending order of their average scores, with earlier predictions used to break any ties. While the original paper used OpenAI’s GPT-3.5-turbo-0613 model for their experiments, for our evaluation, we are using the latest lightweight GPT-4o mini model to perform AutoFL’s experiments.
Table 2. Fault localization accuracy in terms of Top-1, 3, 5, and 10. The numbers in the parenthesis show the number of faults in each project. The result continues in Table 3.
        Project (# faults)         Techniques         Top-1         Top-3         Top-5         Top-10
        Cli (39)         Ochiai         3.0         5.0         10.0         18.0
        DeepFL         11.0         21.0         24.0         28.0
        Grace         14.0         24.0         26.0         30.0
        DepGraph         17.0         24.0         27.0         34.0
        AutoFL         12.0         19.0         19.0         20.0
        LLM4FL         16.0         21.0         23.0         24.0
        Closure (174)         Ochiai         20.0         39.0         70.0         72.0
        DeepFL         46.0         61.0         92.0         99.0
        Grace         51.0         78.0         102.0         121.0
        DepGraph         60.0         99.0         126.0         148.0
        AutoFL         45.0         58.0         65.0         82.0
        LLM4FL         52.0         77.0         102.0         118.0
        Codec (18)         Ochiai         3.0         12.0         17.0         17.0
        DeepFL         5.0         10.0         12.0         16.0
        Grace         6.0         11.0         13.0         17.0
        DepGraph         7.0         10.0         14.0         16.0
        AutoFL         12.0         14.0         14.0         16.0
        LLM4FL         9.0         13.0         13.0         13.0
        Collections (4)         Ochiai         1.0         1.0         2.0         2.0
        DeepFL         1.0         1.0         2.0         2.0
        Grace         1.0         1.0         2.0         2.0
        DepGraph         1.0         2.0         2.0         2.0
        AutoFL         1.0         1.0         1.0         1.0
        LLM4FL         1.0         1.0         1.0         1.0
        Compress (47)         Ochiai         5.0         12.0         17.0         29.0
        DeepFL         22.0         27.0         31.0         38.0
        Grace         23.0         29.0         34.0         42.0
        DepGraph         25.0         33.0         36.0         45.0
        AutoFL         23.0         33.0         34.0         35.0
        LLM4FL         23.0         32.0         34.0         34.0
        Csv (16)         Ochiai         3.0         8.0         10.0         12.0
        DeepFL         7.0         8.0         9.0         11.0
        Grace         6.0         8.0         10.0         12.0
        DepGraph         8.0         9.0         12.0         13.0
        AutoFL         5.0         11.0         12.0         14.0
        LLM4FL         8.0         10.0         10.0         10.0
        Gson (16)         Ochiai         4.0         9.0         9.0         12.0
        DeepFL         8.0         11.0         12.0         12.0
        Grace         11.0         13.0         14.0         15.0
        DepGraph         14.0         15.0         16.0         16.0
        AutoFL         5.0         7.0         10.0         10.0
        LLM4FL         11.0         14.0         14.0         14.0
Table 3. Continued from Table 2. Fault localization accuracy in terms of Top-1, 3, 5, and 10. The numbers in the parenthesis show the number of faults in each project.
        Project (# faults)         Techniques         Top-1         Top-3         Top-5         Top-10
        JacksonCore (26)         Ochiai         6.0         11.0         13.0         14.0
        DeepFL         5.0         5.0         9.0         10.0
        Grace         9.0         13.0         14.0         15.0
        DepGraph         12.0         15.0         15.0         16.0
        AutoFL         10.0         17.0         17.0         17.0
        LLM4FL         12.0         13.0         15.0         15.0
        JacksonXml (6)         Ochiai         0.0         0.0         0.0         0.0
        DeepFL         3.0         3.0         4.0         5.0
        Grace         3.0         3.0         4.0         5.0
        DepGraph         4.0         5.0         5.0         5.0
        AutoFL         2.0         2.0         2.0         3.0
        LLM4FL         4.0         4.0         4.0         4.0
        Jsoup (93)         Ochiai         15.0         40.0         48.0         57.0
        DeepFL         33.0         39.0         46.0         49.0
        Grace         40.0         64.0         72.0         77.0
        DepGraph         53.0         73.0         78.0         83.0
        AutoFL         36.0         52.0         52.0         54.0
        LLM4FL         41.0         56.0         60.0         60.0
        Lang (64)         Ochiai         25.0         45.0         51.0         59.0
        DeepFL         42.0         53.0         55.0         57.0
        Grace         43.0         53.0         57.0         58.0
        DepGraph         48.0         55.0         60.0         61.0
        AutoFL         40.0         57.0         60.0         60.0
        LLM4FL         48.0         55.0         58.0         58.0
        Math (106)         Ochiai         23.0         52.0         62.0         82.0
        DeepFL         52.0         81.0         90.0         95.0
        Grace         64.0         79.0         92.0         97.0
        DepGraph         72.0         92.0         97.0         102.0
        AutoFL         53.0         81.0         87.0         94.0
        LLM4FL         68.0         87.0         92.0         94.0
        Time (26)         Ochiai         6.0         12.0         13.0         16.0
        DeepFL         12.0         15.0         18.0         20.0
        Grace         11.0         16.0         20.0         21.0
        DepGraph         17.0         20.0         21.0         22.0
        AutoFL         13.0         18.0         21.0         22.0
        LLM4FL         14.0         21.0         22.0         23.0
        Mockito (38)         Ochiai         7.0         14.0         18.0         23.0
        DeepFL         10.0         18.0         23.0         26.0
        Grace         16.0         24.0         26.0         29.0
        DepGraph         21.0         29.0         32.0         34.0
        AutoFL         18.0         23.0         29.0         29.0
        LLM4FL         21.0         22.0         26.0         27.0
        Total (675)         Ochiai         121.0         260.0         340.0         413.0
        DeepFL         257.0         353.0         427.0         468.0
        Grace         298.0         416.0         486.0         541.0
        DepGraph         359.0         481.0         541.0         597.0
        AutoFL         275.0         393.0         423.0         457.0
        LLM4FL         328.0         426.0         474.0         495.0
Results. LLM4FL outperforms the LLM-based baseline, AutoFL, by achieving a much higher Pass@1 (328 v.s. 275). Table 2 and 3 show the fault localization results of LLM4FL and the baseline techniques. Between the two LLM-based techniques, LLM4FL achieves a better Top@N than AutoFL when N is 1, 3, 5, and 10. In the Top-1 metric, LLM4FL locates the correct fault in 328 cases, compared to AutoFL’s 275, representing a 19.27% improvement. Similarly, in the Top-3 and Top-5 metrics, LLM4FL achieves scores of 426 and 474, respectively, outperforming AutoFL’s results of 393 and 423. Even in the broader Top-10 metric, LLM4FL shows an 8.32% advantage, with a score of 495 compared to AutoFL’s 457. These numbers highlight LLM4FL’s enhanced ability to pinpoint faulty methods more accurately and efficiently, reinforcing its superiority over AutoFL for fault localization tasks.
LLM4FL shows higher Top-1 and Top-3 compared to all other non-LLM-based techniques, except for DepGraph. For the Top-1 metric, LLM4FL scores 328, which is 171.07% higher than Ochai’s score of 121, 27.63% higher than DeepFL’s score of 257, and 10.07% better than Grace’s result of 298. One exception is DepGraph, which achieves a Top-1 of 359, 8.64% higher than LLM4FL. As the range expands to Top-3 and beyond, LLM4FL continues to demonstrate its robustness, significantly outperforming DeepFL and maintaining strong performance alongside Grace. It is important to note that DeepFL, Grace, and DepGraph are supervised methods trained and evaluated using leave-one-out cross-validation, following the original study protocols. Despite LLM4FL being a zero-shot approach without task-specific training, its ability to perform competitively with these supervised techniques further highlights its effectiveness. The fact that LLM4FL achieves strong results without the need for extensive training data or cross-validation underscores its potential and sheds light on future research directions, especially when training models on specific data sets may not always be feasible.
LLM4FL outperforms the other LLM-based technique, AutoFL, by achieving a 19.27% higher Top-1 and delivering competitive results compared to supervised techniques like DeepFL and Grace, even without task-specific training. The findings highlight the potential of using LLM4FL for zero-shot fault localization.
4.2. RQ2: How do different components affect LLM4FL’s fault localization accuracy?
Motivation. In this research question, we study the influence of each individual component within LLM4FL on its fault localization performance. LLM4FL employs a combination of advanced techniques, including coverage splitting to handle token limitations, prompt chaining for fault navigation, self-reflection for iterative refinement of the agents’ decisions, and chain-of-thought reasoning for re-ranking suspicious methods. Each of these components plays a distinct role in the overall process. To study their effects, we investigate how the removal of any one of these components affects the LLM4FL’s ability to accurately localize faults. By isolating the effects of each technique, we can better understand which components are critical to maintaining or improving the overall accuracy of the LLM4FL and provide insights to improving future LLM-based fault localization techniques.
Approach. To evaluate the impact of each component, we designed four experimental configurations: LLM4FLw/o CodeNavigation, LLM4FLw/o Divide&Conquer, LLM4FLw/o Reflection, and LLM4FLw/o CoT. For each configuration, one component is systematically removed, and we compare the result with the full-featured baseline of LLM4FL. This allows us to directly measure how much each component contributes to the fault localization process.
LLM4FLw/o CodeNavigation removes the prompt chaining mechanism, which essentially means the agents no longer collaborate do fault navigation through multiple rounds of reasoning. Instead, the LLM4FL uses a single prompt to perform fault localization without iterative agent communication and fault navigation. This configuration tests whether the multi-step agent collaboration improves the ranking and selection of faulty methods or if a single round of communication is sufficient.
LLM4FLw/o Divide&Conquer removes the grouping of the covered methods, giving the agents the entire coverage data at once instead of dividing it into smaller, manageable groups. Coverage segmentation addresses token limitations in LLMs, so removing it explores the impact of feeding the full dataset to the agents in one step. We aim to see how handling large amounts of data in a single input influences the fault localization result, as it may overwhelm the model or reduce precision.
LLM4FLw/o Reflection removes the self-reflection technique, which is used to allow agents to review and refine their initial candidate selection and ranking. Without this self-reflection step, the agents rely solely on their initial assessments without iterative improvements.
LLM4FLw/o CoT disables the chain-of-thought reasoning process, which asks the agents to think deeper by asking LLMs to generate potential fixes and re-rank the suspicious methods. The chain-of-thought approach is designed to enhance logical reasoning and ensure that intermediate results are critically evaluated and refined.
Table 4. Impacts of removing different components in LLM4FL on Top-1, 3, 5, and 10. The numbers in the parenthesis show the percentage changes compared to LLM4FL with all the components.
       Project (# faults)        Techniques        Top-1        Top-3        Top-5        Top-10
       Cli (39)        LLM4FL        16        21        23        24
       LLM4FLw/o CodeNavigation        11 (-31.25%)        19 (-9.52%)        21 (-8.7%)        21 (-12.5%)
       LLM4FLw/o Divide&Conquer        11 (-31.25%)        17 (-19.05%)        17 (-26.09%)        19 (-20.83%)
       LLM4FLw/o Reflection        11 (-31.25%)        19 (-9.52%)        21 (-8.7%)        21 (-12.5%)
       LLM4FLw/o CoT        16 (0.0%)        21 (0.0%)        23 (0.0%)        24 (0.0%)
       Closure (174)        LLM4FL        52        77        102        118
       LLM4FLw/o CodeNavigation        32 (-38.46%)        45 (-41.56%)        53 (-48.04%)        53 (-55.08%)
       LLM4FLw/o Divide&Conquer        27 (-48.08%)        40 (-48.05%)        49 (-51.96%)        50 (-57.63%)
       LLM4FLw/o Reflection        50 (-3.85%)        74 (-3.9%)        92 (-9.8%)        109 (-7.63%)
       LLM4FLw/o CoT        52 (0.0%)        77 (0.0%)        102 (0.0%)        118 (0.0%)
       Codec (18)        LLM4FL        9        13        13        13
       LLM4FLw/o CodeNavigation        8 (-11.11%)        12 (-7.69%)        13 (0.0%)        13 (0.0%)
       LLM4FLw/o Divide&Conquer        8 (-11.11%)        11 (-15.38%)        12 (-7.69%)        13 (0.0%)
       LLM4FLw/o Reflection        9 (0.0%)        12 (-7.69%)        13 (0.0%)        13 (0.0%)
       LLM4FLw/o CoT        9 (0.0%)        13 (0.0%)        13 (0.0%)        13 (0.0%)
       Compress (47)        LLM4FL        23        32        34        34
       LLM4FLw/o CodeNavigation        23 (0.0%)        32 (0.0%)        34 (0.0%)        34 (0.0%)
       LLM4FLw/o Divide&Conquer        23 (0.0%)        28 (-12.5%)        30 (-11.76%)        31 (-8.82%)
       LLM4FLw/o Reflection        23 (0.0%)        32 (0.0%)        35 (2.94%)        35 (2.94%)
       LLM4FLw/o CoT        23 (0.0%)        32 (0.0%)        34 (0.0%)        34 (0.0%)
       Csv (16)        LLM4FL        8        10        10        10
       LLM4FLw/o CodeNavigation        8 (0.0%)        8 (-20.0%)        9 (-10.0%)        9 (-10.0%)
       LLM4FLw/o Divide&Conquer        8 (0.0%)        8 (-20.0%)        9 (-10.0%)        9 (-10.0%)
       LLM4FLw/o Reflection        7 (-12.5%)        10 (0.0%)        10 (0.0%)        10 (0.0%)
       LLM4FLw/o CoT        8 (0.0%)        9 (-10.0%)        10 (0.0%)        10 (0.0%)
       Gson (16)        LLM4FL        11        14        14        14
       LLM4FLw/o CodeNavigation        9 (-18.18%)        15 (7.14%)        15 (7.14%)        15 (7.14%)
       LLM4FLw/o Divide&Conquer        10 (-9.09%)        14 (0.0%)        15 (7.14%)        15 (7.14%)
       LLM4FLw/o Reflection        11 (0.0%)        12 (-14.29%)        13 (-7.14%)        14 (0.0%)
       LLM4FLw/o CoT        11 (0.0%)        14 (0.0%)        14 (0.0%)        14 (0.0%)
       JacksonCore (26)        LLM4FL        12        13        15        15
       LLM4FLw/o CodeNavigation        9 (-25.0%)        13 (0.0%)        14 (-6.67%)        14 (-6.67%)
       LLM4FLw/o Divide&Conquer        9 (-25.0%)        13 (0.0%)        13 (-13.33%)        13 (-13.33%)
       LLM4FLw/o Reflection        10 (-16.67%)        13 (0.0%)        13 (-13.33%)        13 (-13.33%)
       LLM4FLw/o CoT        12 (0.0%)        13 (0.0%)        15 (0.0%)        15 (0.0%)
       JacksonXml (6)        LLM4FL        4        4        4        4
       LLM4FLw/o CodeNavigation        2 (-50.0%)        3 (-25.0%)        3 (-25.0%)        3 (-25.0%)
       LLM4FLw/o Divide&Conquer        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)
       LLM4FLw/o Reflection        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)
       LLM4FLw/o CoT        3 (-25.0%)        4 (0.0%)        4 (0.0%)        4 (0.0%)
       Jsoup (93)        LLM4FL        41        56        60        60
       LLM4FLw/o CodeNavigation        40 (-2.44%)        55 (-1.79%)        58 (-3.33%)        58 (-3.33%)
       LLM4FLw/o Divide&Conquer        36 (-12.2%)        50 (-10.71%)        51 (-15.0%)        51 (-15.0%)
       LLM4FLw/o Reflection        38 (-7.32%)        53 (-5.36%)        54 (-10.0%)        54 (-10.0%)
       LLM4FLw/o CoT        40 (-2.44%)        55 (-1.79%)        59 (-1.67%)        60 (0.0%)
       Lang (64)        LLM4FL        48        55        58        58
       LLM4FLw/o CodeNavigation        42 (-12.5%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
       LLM4FLw/o Divide&Conquer        43 (-10.42%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
       LLM4FLw/o Reflection        42 (-12.5%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
       LLM4FLw/o CoT        48 (0.0%)        55 (0.0%)        59 (1.72%)        59 (1.72%)
       Math (106)        LLM4FL        68        87        92        94
       LLM4FLw/o CodeNavigation        59 (-13.24%)        79 (-9.2%)        85 (-7.61%)        85 (-9.57%)
       LLM4FLw/o Divide&Conquer        49 (-27.94%)        68 (-21.84%)        70 (-23.91%)        78 (-17.02%)
       LLM4FLw/o Reflection        57 (-16.18%)        80 (-8.05%)        85 (-7.61%)        85 (-9.57%)
       LLM4FLw/o CoT        68 (0.0%)        87 (0.0%)        91 (-1.09%)        94 (0.0%)
       Mockito (38)        LLM4FL        21        22        26        27
       LLM4FLw/o CodeNavigation        18 (-14.29%)        23 (4.55%)        23 (-11.54%)        23 (-14.81%)
       LLM4FLw/o Divide&Conquer        15 (-28.57%)        20 (-9.09%)        21 (-19.23%)        21 (-22.22%)
       LLM4FLw/o Reflection        18 (-14.29%)        20 (-9.09%)        20 (-23.08%)        22 (-18.52%)
       LLM4FLw/o CoT        21 (0.0%)        22 (0.0%)        26 (0.0%)        27 (0.0%)
       Time (26)        LLM4FL        14        21        22        23
       LLM4FLw/o CodeNavigation        12 (-14.29%)        20 (-4.76%)        22 (0.0%)        22 (-4.35%)
       LLM4FLw/o Divide&Conquer        10 (-28.57%)        16 (-23.81%)        17 (-22.73%)        20 (-13.04%)
       LLM4FLw/o Reflection        12 (-14.29%)        19 (-9.52%)        19 (-13.64%)        22 (-4.35%)
       LLM4FLw/o CoT        14 (0.0%)        22 (4.76%)        23 (4.55%)        23 (0.0%)
       Total (675)        LLM4FL        327        425        473        494
       LLM4FLw/o CodeNavigation        273 (-16.51%)        378 (-11.06%)        409 (-13.53%)        409 (-17.21%)
       LLM4FLw/o Divide&Conquer        251 (-23.24%)        341 (-19.76%)        365 (-22.83%)        381 (-22.87%)
       LLM4FLw/o Reflection        290 (-11.31%)        400 (-5.88%)        436 (-7.82%)        459 (-7.09%)
       LLM4FLw/o CoT        325 (-0.61%)        424 (-0.24%)        473 (0.0%)        495 (0.2%)
Results. While all components help improve the results, including coverage grouping and prompt chaining provide the largest improvement to fault localization results (23% and 17% in Top-1). Table 4 shows the Top-1, 3, 5, and 10 scores when each component is removed. Removing coverage splitting has the largest overall impact across all scores, reducing Top-N by 19% to 23%. Removing prompt chaining has the second largest impact (11% to 17%). At the individual project level, these two components also have the largest impact in Top-1 in 9/13 studied projects. Our finding shows that employing sorted coverage grouping following the divide and conquer technique and agent communication significantly improves fault localization results. Future research should consider these techniques when designing fault localization techniques.
Although there is no oracle during the fault localization process, asking LLMs to self-reflection still helps improve the overall Top-1 by 11%. LLMs often suffer from hallucinations, especially when there is a lack of feedback from external oracles (Xu et al., 2024; Huang et al., 2023). Even though we did not provide any groundtruth or external feedback to LLM, we find that self-reflection is still effective in improving fault localization results. Self-reflect brings 6% to 11% improvement across the Top-N metrics. Our finding suggests that future studies should consider self-reflection even if there is no external feedback. Interestingly, we did not see large differences after removing chain-of-thought reasoning via test generation. In some cases, removing it even slightly improves the fault localization results. One reason may be that LLM4FL already includes self-reflection to make LLM think deeply about the result, and adding a chain-of-thought did not further improve the result. Our finding highlights the effectiveness of self-reflection, which should be considered in future fault localization results.
The results show that each component of LLM4FL contributes to its overall fault localization performance, with coverage splitting and prompt chaining having the largest positive impact. Removing these components leads to significant declines in accuracy, confirming their crucial role in handling token limitations and enabling more effective multi-agent collaboration.
4.3. RQ3: Does order matter in the initial list of methods provided to the LLM?
Motivation. LLM4FL divides the coverage data into segments to address the token size limitation of LLMs. We sort the methods using the Ochiai scores before segmentation, though different sorting mechanisms may affect the final fault localization result. Although LLM4FL eventually visits and assesses every method, a recent study (Chen et al., 2024) observes that the order of premises affects LLM’s results. However, whether this effect extends to software engineering tasks, particularly fault localization, remains unclear. Hence, in this research question, we investigate whether the order of methods within the segments affects the LLM’s fault localization performance.
Approach. To test the effect of method ordering, we experiment with three distinct sorting strategies: LLM4FLExecution, LLM4FLOchiai (the default sorting in LLM4FL), and LLM4FLDepGraph. Each strategy provides a different way to sort the methods before they are segmented for further analysis.
LLM4FLExecution: In this baseline approach, we use the unsorted list of methods executed during testing, as generated by Gzoltar (Campos et al., 2012). This default list represents the natural execution order of the methods, with no explicit ranking or prioritization. By providing the LLM with methods based on the execution sequence, we establish a control case to measure its performance without any ranking influence. The result shows how the order of premises in reasoning tasks can lead to varied LLM performance in software engineering context (Chen et al., 2024).
LLM4FLOchiai: As discussed in Section 3.1, we apply Ochiai to sort the methods during the segmentation process. Ochiai is unsupervised and is efficient to compute. We hypothesize that providing the LLM with methods sorted by their suspiciousness score will lead to more effective fault localization, as the model will focus on the most likely faulty candidates earlier in the process.
LLM4FLDepGraph: This approach uses the ranking produced by DepGraph, a state-of-the-art Graph Neural Network (GNN)-based fault localization technique (Li et al., 2015; Rafi et al., 2024), to sort the methods during the segmentation process. DepGraph ranks methods based on structural code dependencies and code change history. As shown in RQ1, DepGraph shows the highest fault localization accuracy among all the techniques, surpassing LLM4FLOchiai. By examining the fault localization result after sorting the methods using DepGraph’s scores, we can better study if the initial order affects LLM’s results, even though LLM eventually visits all the methods.
Figure 4. Fault localization results when using different method sorting strategies during the segmentation process.
Results. Method ordering has a significant impact on LLM’s fault localization result, with up to 22% difference in Top-1 (from 299 to 366). Figure 4 shows the fault localization results when using different sorting strategies. When methods were presented in an execution-based order, LLM4FLExecution achieved a Top-1 score of 299, gradually increasing to 402 for Top-3, 431 for Top-5, and reaching 462 for Top-10. This performance establishes a baseline, showing how the LLM behaves without strategic ordering. However, sorting methods with the lightweight Ochiai scores resulted in noticeable improvements. LLM4FLOchiai improved the Top-1 score to 323, an 8% increase over LLM4FLExecution. The improvements also extended to other metrics, with the Top-3 score increasing to 421, the Top-5 reaching 469, and the Top-10 rising to 490.
LLM4FLDepGraph provides further improvement to the already-promising result of DepGraph, indicating method ordering is critical to LLM4FL, or LLM-based fault localization in general. LLM4FLDepGraph achieved the highest Top-1 score of 366, which is significantly outperforming both LLM4FLExecution and LLM4FLOchiai. The improvement was consistent across all the metrics. For Top-3, the score rose to 492, providing a substantial boost over LLM4FLOchiai. Similarly, LLM4FLDepGraph excelled in the Top-5 and Top-10 categories, reaching 543 and 548, respectively. We also see that LLM4FLDepGraph has better Top-1, 3, and 5 scores compared to DepGraph. This consistent improvement underscores the importance of method ordering in enhancing the accuracy of LLM-based fault localization. This underscores the importance of method ordering for LLM-based fault localization, showing that lightweight techniques like Ochiai can achieve substantial performance gains without the heavy computational burden. Striking this balance between accuracy and efficiency makes these approaches particularly suitable for a wide range of fault localization tasks. However, it is important to recognize that LLM4FLDepGraph is computationally expensive due to its reliance on graph neural networks (GNNs). While LLM4FLDepGraph delivers top-tier results, LLM4FLOchiai also offers a strong, efficient alternative, delivering good localization accuracy at a fraction of the computational cost due to the unsupervised nature of Ochiai. In other words, LLM4FLOchiai may be more easily adapted in practice.
Nevertheless, our finding establishes a new research direction for LLM-based fault localization. It demonstrates that intelligent method ordering strategies significantly impact the result of LLM-based fault localization. This approach opens up further opportunities for optimizing LLM-based fault localization by exploring more advanced ordering techniques and how different premises of ordering affect other software engineering tasks.
The findings highlight that method ordering plays a crucial role in improving the performance of LLM-based fault localization, with a difference of up to 22% in Top-1 scores. While LLM4FLDepGraph delivers the best results, the lightweight LLM4FLOchiai offers a highly efficient alternative, providing significant accuracy gains with far lower computational costs, making it more practical for real-world adoption.
5. Threats To Validity
Internal Validity. A potential threat to internal validity is the risk of data leakage in large language models (LLMs), where the model might have been exposed to data similar to the benchmark or specific fault localization cases during training. This could result in artificially inflated performance as the model may have prior knowledge of the evaluation data. Nevertheless, we mitigate this risk by preventing any content from being entered into ChatGPT that is related to the project name, the human-written bug report, or the bug ID.
External Validity. The primary threat to external validity is the generalizability of our results. Our evaluation is based on Defects4J, a well-established dataset in the software engineering community. Although this dataset includes real-world bugs, the systems studied are primarily Java-based, and it remains uncertain whether our findings will generalize to other programming languages or domains.
Construct Validity. Construct validity relates to whether the metrics we used accurately measure the performance of fault localization techniques. We used widely accepted Top-N metrics, which are commonly utilized in prior fault localization studies. However, our results are based on the assumption that developers primarily focus on the top-ranked faulty methods. Although this assumption aligns with previous research, different development practices could influence the effectiveness of our approach.
6. Conclusion
In this paper, we introduced LLM4FL, an LLM-agent-based fault localization technique that addresses the challenges of traditional fault localization methods and the limitations of current LLM-based approaches. By combining SBFL rankings with a divide-and-conquer strategy, LLM4FL groups large coverage data into manageable pieces, enabling effective analysis within the token limitations of LLMs. The use of multiple agents, prompt chaining, and self-reflection techniques allows for iterative refinement of fault localization, while chain-of-thought reasoning further enhances accuracy by generating potential fixes and re-rank the methods. Our evaluation on the Defects4J (V2.0.0) benchmark demonstrated that LLM4FL outperforms existing LLM-based techniques, achieving 19.27% higher Top-1 accuracy compared to AutoFL, and surpasses supervised approaches such as DeepFL and Grace without task-specific training. We also found that key components like coverage splitting and prompt chaining are essential to LLM4FL’s success, with method ordering playing a significant role in performance, improving Top-1 accuracy by up to 22%. Overall, LLM4FL presents a scalable and efficient solution for fault localization, providing improved accuracy and reducing computational costs, making it practical for real-world software projects. Future work will explore expanding LLM4FL’s capabilities for larger and more diverse codebases, further refining the agent collaboration and reasoning mechanisms.

### v2

@@
-Enhancing Fault Localization Through Ordered Code Analysis with LLM Agents and Self-Reflection
+A Multi-Agent Approach to Fault Localization via Graph-Based Retrieval and Reflexion
 Abstract.
-Locating and fixing software faults is a time-consuming and resource-intensive task in software development.
-Traditional fault localization methods, such as Spectrum-Based Fault Localization (SBFL), rely on statistical analysis of test coverage data but often suffer from lower accuracy.
-Learning-based techniques, while more effective, require extensive training data and can be computationally expensive.
-Recent advancements in Large Language Models (LLMs) offer promising improvements in fault localization by enhancing code comprehension and reasoning.
-However, these LLM-based techniques still face challenges, including token limitations, degraded performance with long inputs, and difficulties managing large-scale projects with complex systems involving multiple interacting components.
-To address these issues, we introduce LLM4FL, a novel LLM-agent-based fault localization approach that integrates SBFL rankings with a divide-and-conquer strategy.
-By dividing large coverage data into manageable groups and employing multiple LLM agents through prompt chaining, LLM4FL navigates the codebase and localizes faults more effectively.
-The approach also incorporates self-reflection and chain-of-thought reasoning, enabling agents to iteratively generate fixes and re-rank suspicious methods.
-We evaluated LLM4FL on the Defects4J (V2.0.0) benchmark, comprising 675 real-world faults from 14 open-source Java projects.
-Our results demonstrate that LLM4FL outperforms AutoFL by 19.27% in Top-1 accuracy and surpasses state-of-the-art supervised techniques such as DeepFL and Grace, all without task-specific training.
-Additionally, we highlight the impact of coverage splitting and prompt chaining on fault localization performance and show that different method ordering can improve Top-1 accuracy by up to 22%.
+Identifying and resolving software faults remains a challenging and resource-intensive process.
+Traditional fault localization techniques, such as Spectrum-Based Fault Localization (SBFL), leverage statistical analysis of test coverage but often suffer from limited accuracy.
+While learning-based approaches improve fault localization, they demand extensive training datasets and high computational resources.
+Recent advances in Large Language Models (LLMs) offer new opportunities by enhancing code understanding and reasoning.
+However, existing LLM-based fault localization techniques face significant challenges, including token limitations, performance degradation with long inputs, and scalability issues in complex software systems.
+To overcome these obstacles, we propose LLM4FL, a multi-agent fault localization framework that utilizes three specialized LLM agents.
+First, the Context Extraction Agent applies an order-sensitive segmentation strategy to partition large coverage data within the LLM’s token limit, analyze failure context, and prioritize failure-related methods.
+The extracted data is then processed by the Debugger Agent, which employs graph-based retrieval-augmented code navigation to reason about failure causes and rank suspicious methods.
+Finally, the Reviewer Agent re-evaluates the identified faulty methods using verbal reinforcement learning, engaging in self-criticism and iterative refinement.
+Evaluated on the Defects4J (V2.0.0) benchmark, which includes 675 faults from 14 Java projects, LLM4FL achieves an 18.55% improvement in Top-1 accuracy over AutoFL and 4.82% over SoapFL.
+It outperforms supervised techniques such as DeepFL and Grace, all without requiring task-specific training.
+Furthermore, its coverage segmentation and prompt chaining strategies enhance performance, increasing Top-1 accuracy by up to 22%.
+Fault localization, large language model, agent
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
 1. Introduction
@@
 Research shows that software development teams allocate more than half of their budgets to testing and debugging activities (Hait and Tassey, 2002; Alaboudi and LaToza, 2021).
-As software systems become increasingly complex, the demand for more efficient and accurate fault localization techniques continues to grow.
+As software systems become increasingly complex, the demand for more accurate fault localization techniques grows.
 Hence, to assist developers and reduce debugging costs, researchers have developed various fault localization techniques (Li et al., 2019; Lou et al., 2021; Qian et al., 2023b; Li et al., 2021; Abreu et al., 2009; Sohn and Yoo, 2017).
-These techniques analyze code coverage and program execution to identify the most likely faulty components, assisting developers in finding the fault.
-However, despite the advances in fault localization techniques, many existing approaches still struggle with scalability and precision.
+These techniques analyze code coverage and program execution to identify the most likely faulty code, assisting developers in finding the fault.
+Despite the advances in fault localization, many existing techniques still struggle with scalability and precision.
 Traditional methods, such as Spectrum-Based Fault Localization (SBFL), use statistical analysis to analyze coverage data from passing and failing test cases to rank suspicious code elements (Abreu et al., 2006).
@@
 Their reliance on statistical correlations between test failures and code coverage does not always capture the deeper semantic relationships needed for more accurate fault localization (Wong et al., 2016; Xie et al., 2013; Le et al., 2013).
-To address these issues, recent techniques applied machine learning and deep learning models to improve fault localization (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021, 2019).
+Recent techniques applied machine learning (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021, 2019) and deep learning models (Lou et al., 2021; Qian et al., 2023b; Rafi et al., 2024; Lou et al., 2021) to address these issues to improve fault localization.
 These methods enhance the ranking of suspicious code elements by incorporating additional information like code complexity, text similarity, and historical fault data.
-Researchers also leveraged models like Graph Neural Networks (GNNs) to represent code structures and achieved state-of-the-art fault localization accuracy (Lou et al., 2021; Qian et al., 2023b; Rafi et al., 2024; Lou et al., 2021).
-However, these techniques often require extensive training data and significant training time.
-Recent advances in Large Language Models (LLMs) have shown great potential for fault localization due to their strong language comprehension and generation capabilities (Abedu et al., 2024; Lin et al., 2024).
-LLMs trained on extensive programming datasets can understand code structure, interpret error messages, and even suggest fixes for common software bugs (Kang et al., 2024; Wu et al., 2023; Pu et al., 2023; Li et al., 2023).
+However, these techniques often require extensive training data that may not be available.
+Recent advances in Large Language Models (LLMs) have shown great potential for software fault localization due to their strong language comprehension and generation capabilities (Abedu et al., 2024; Lin et al., 2024b).
+LLMs trained on extensive programming datasets can understand code structure, interpret error messages, and even suggest bug fixes (Wu et al., 2023; Pu et al., 2023; Li et al., 2023).
 These models, with their ability to analyze and process both natural language and code, present an opportunity to significantly improve traditional fault localization methods by incorporating deeper semantic analysis and context-aware reasoning.
-Wu et al.
-(Wu et al., 2023), directly present LLMs with faulty methods or classes with test failure information and ask to locate the issue, which provides valuable insights.
-Kang et al.
-(Kang et al., 2024) which operates as an automated fault localization technique that leverages LLMs to localize fault given a single failing test.
-It focuses on method-level fault localization and provides bug location and also a natural language explanation of why a particular code location is likely to be faulty.
-Despite their potential, existing LLM-based fault localization techniques face several challenges.
-The token limitations of LLMs restrict their ability to effectively process long code files or large sets of code coverage data, which is often required when dealing with large-scale software systems (Hadi et al., 2023; Hou et al., 2023; Wu et al., 2023).
-Additionally, the performance of LLMs can degrade when applied to complex systems that require the model to reason over multiple interacting components, making it difficult to maintain accuracy and consistency across broader projects (Levy et al., 2024; Liu et al., 2024).
-Furthermore, current LLM-based techniques have yet to fully explore how these models can be effectively integrated with traditional fault localization techniques to maximize their strengths in a complementary and efficient manner (Wu et al., 2023; Kang et al., 2024; Yang et al., 2024; Bin Murtaza et al., 2024).
-In this paper, we propose LLM4FL, an LLM-based fault localization.
-To address the challenges of analyzing large-scale software projects, where code coverage and complexity often exceed LLM token limits (Wu et al., 2023; Liu et al., 2024), LLM4FL implements a divide-and-conquer strategy.
-Before applying the divide-and-conquer strategy, we use an SBFL technique to sort the covered methods, building on findings from prior research (Chen et al., 2024) that demonstrate LLM performance improves when the order of instructions is carefully considered.
-Then, we divide the coverage data into manageable groups that the LLM can process within its token limits.
-Note that, regardless of the order, LLM4FL eventually analyzes every covered method.
-In addition to using a divide-and-conquer strategy, LLM4FL takes inspiration from how human developers debug software.
-Developers typically analyze multiple types of information, including error messages, stack traces, and code snippets, to incrementally narrow down potential faulty components (Alaboudi and LaToza, 2021; Böhme et al., 2017).
-LLM4FL emulates this process by utilizing two LLM agents collaborating to iteratively and autonomously navigate the code to locate the faults.
-LLM4FL implements a Tester and a Debugger Agent, each tasked with specialized tools to assist in the fault localization process.
-The Tester Agent identifies and prioritizes suspicious methods by analyzing the failing test, and stack traces in different groups of covered methods to list a list of highly suspicious methods.
-It mimics the developer’s process of investigating suspicious areas in the code by understanding the context of error messages.
-Meanwhile, the Debugger Agent thoroughly evaluates the given list of candidate methods by navigating the code and ranks them based on its analysis.
-These two agents communicate through a prompt chaining mechanism that allows them to share insights and build upon each other’s findings.
+However, there are several challenges in using LLMs for fault localization.
+Firstly, LLMs face significant difficulties when analyzing large-scale repository-level data, primarily due to inherent token limitations, as highlighted in recent studies (Hadi et al., 2023; Hou et al., 2023; Wu et al., 2023).
+Secondly, the performance of LLMs tends to decrease when they are applied to complex systems that require reasoning about multiple interacting components.
+This complexity can negatively impact the performance of LLMs (Liu et al., 2024; Wu et al., 2023).
+LLMs may struggle to understand the inter-procedural method call relationships, an essential aspect for understanding how faults propagate across the entire software repository (Chatterjee et al., 2015; Levy et al., 2024; Liu et al., 2024).
+In this paper, we propose LLM4FL, a multi-agent LLM-based fault localization technique designed to analyze large-scale software projects at the repository level.
+LLM4FL leverages a multi-agent framework that emulates developers’ fault localization process.
+LLM4FL employs three LLM agents, each with a specialized role to enhance fault localization effectiveness.
+The Context Extraction Agent divides the methods covered by the failing test cases into smaller groups, each within LLM’s token limitation.
+Then, the agent analyzes the test code and test failure messages to reason about possible failure causes.
+This helps prioritize the most suspicious covered methods from each group before passing them on to the next agent.
+Next, the Debugger Agent analyzes each covered method and performs graph-based, retrieval-augmented (Graph-RAG) code navigation.
+It traverses the prioritized methods’ code iteratively along call dependencies to analyze and rank suspicious methods.
+Finally, the Reviewer Agent re-ranks the identified fault methods collected from all groups using verbal reinforcement learning (Shinn et al., 2024), criticizing and refining each response to ensure more precise fault localization.
 We evaluated LLM4FL using the Defects4J (V2.0.0) benchmark (Just et al., 2014), which contains 675 real-world faults from 14 open-source Java projects.
-Our results demonstrate that LLM4FL surpasses LLM-based technique AutoFL (Kang et al., 2024), by achieving 19.27% higher Top-1 accuracy.
+Our results demonstrate that LLM4FL surpasses LLM-based technique AutoFL (Kang et al., 2024) and AgentFL/SoapFL (Qin et al., 2024) by achieving 18.55% and 4.82% higher Top-1 accuracy, respectively.
+LLM4FL is also the cheapest among the three, costing only \$0.05 per fault.
 Additionally, LLM4FL outperforms supervised techniques such as DeepFL (Li et al., 2019) and Grace (Lou et al., 2021), even without task-specific training.
 We also analyzed the impact of individual components within LLM4FL on fault localization accuracy.
-Our findings indicate that each component plays a significant role in its performance, with coverage splitting and prompt chaining contributing the most.
-When these components are removed, accuracy drops considerably, underscoring their importance in managing token limitations and facilitating efficient multi-agent collaboration.
+Our findings indicate that each component plays a significant role in performance.
+Among these, dividing the covered methods into smaller groups, and Graph-RAG-based code navigation contributes the most.
 Moreover, we examined whether the initial ordering of methods provided to the LLM influences performance.
-The results reveal that method ordering is important, with a Top-1 accuracy difference of up to 22% when comparing an execution-based ordering and the order provided by DepGraph (Rafi et al., 2024).
+The results reveal that method ordering is important.
+Even though LLM eventually visits all methods, the Top-1 accuracy still has a difference of up to 22% when comparing an execution-based ordering and the order provided by DepGraph (Rafi et al., 2024).
 The paper makes the following contributions:
@@
 This technique groups large coverage data and ranks the covered methods using an SBFL formula.
-By utilizing multiple agents and prompt chaining, LLM4FL navigates the code iteratively to effectively identify and localize faults.
-  • LLM4FL demonstrates superior performance, surpassing AutoFL (Kang et al., 2024) by 19.27% in Top-1 accuracy.
-It also outperforms state-of-the-art supervised techniques like DeepFL and Grace, achieving these results without requiring task-specific training.
-  • Our analysis of LLM4FL’s components shows that key features like coverage splitting and prompt chaining are essential to its fault localization accuracy.
-Removing these features leads to significant performance declines, emphasizing their importance in handling token limitations and enabling effective agent collaboration.
-  • We further investigate the effect of method ordering on LLM performance.
-The study reveals that different ordering can enhance fault localization accuracy by up to 22% in Top-1 scores.
-While LLM4FLDepGraph achieves the highest overall accuracy, LLM4FLOchiai offers a more efficient solution, balancing accuracy gains with lower computational overhead, making it practical for broader use.
-In short, we provide a strategy to mitigate the token limitation issues in LLM-based fault localization and highlight the impact of initial method ordering.
-The findings may help inspire future research to improve LLM-based fault localization for large-scale software projects.
+Using multiple agents and code navigation based on Graph-RAG, LLM4FL analyzes the repository iteratively to identify and localize faults.
+  • We conducted an extensive evaluation and LLM4FL demonstrates superior performance, surpassing AutoFL (Kang et al., 2024) by 18.55% and AgentFL/SoapFL (Qin et al., 2024) by 4.82% in Top-1 accuracy.
+It also outperforms supervised techniques like DeepFL and Grace, achieving these results without requiring task-specific training.
+  • Our analysis of LLM4FL’s components shows that key components like dividing methods into smaller groups and code navigation are essential to its fault localization accuracy.
+Removing these features leads to performance declines, emphasizing their importance in handling token limitations and effective fault analysis.
+  • We find that the different ordering on the initial method list passed to LLM can affect fault localization accuracy by up to 22% in Top-1 scores, even though LLM eventually visits all the methods.
+In short, we provide a strategy to mitigate the token limitation issues and analyze repository-level data in LLM-based fault localization.
+We also highlight the impact of initial method ordering for LLM’s input.
+The findings may help inspire future research on LLM-based fault localization for large-scale software projects.
 Paper Organization.
-Section 2 discusses related work.
+Section 2 discusses background and related work.
 Section 3 describes our technique, LLM4FL.
@@
 Section 6 concludes the paper.
+Table 1. Comparison of LLM-based fault localization techniques.
+Aspect LLM4FL (this work) AutoFL (Kang et al., 2024) SoapFL/AgentFL (Qin et al., 2024) AutoCodeRover (Zhang et al., 2024)
+Navigating large code repository The agent autonomously navigates the repository, iteratively following an inter-procedural call graph to retrieve and explore connected methods.
+Retrieve class and method signatures related to the failure, then let the LLM decide which methods need deeper analysis.
+Analyze the test failures to identify the most suspicious classes, then iteratively analyze related methods.
+Start with the file or method names extracted from issue reports, then iteratively search the repository based on called methods in the retrieved source code.
+Handling token limitation Divides methods into groups and analyzes each group separately.
+Provide only class and method signatures to the LLM, then let it decide which methods need deeper analysis.
+Use LLMs to generate description for classes and methods using their signatures.
+Then, the LLM decides which classes/methods are relevant to the test failures for further analysis.
+Iteratively search for relevant classes and methods based on the issue description and prioritize methods ranked by SBFL techniques.
+Ranking of faulty methods Hierarchical ranking.
+First, it finds the suspicious methods and ranks them in each group.
+Then, it employs verbal reinforcement learning to combine and re-rank the suspicious methods from all groups.
+Re-run the fault localization process multiple times and do a majority vote on the result.
+Utilize a multi-round dialogue strategy that allows LLM agents to discuss the code to score and rank faulty methods.
+Does not rank methods; it selects a method as faulty from the retrieved context and generates a patch.
 2. Background and Related Work
@@
 The training involves self-supervised learning objectives that enable these models to develop a deep understanding of language and generate contextually relevant and semantically coherent text.
-LLMs have shown substantial capability in tasks that involve complex language comprehension and generation (Abedu et al., 2024; Lin et al., 2024), such as code recognition and generation.
+LLMs have shown substantial capability in tasks that involve complex language comprehension and generation (Abedu et al., 2024; Lin et al., 2024b), such as code recognition and generation.
 Recent research has leveraged LLMs in software engineering tasks, particularly in fault localization (Kang et al., 2024; Qin et al., 2024; Yang et al., 2024), where they assist in identifying the faulty code groups responsible for software errors.
@@
 They leverage functionalities like memory management (Zhou et al., 2023) and tool integration (Xia et al., 2024; Roy et al., 2024) to handle multi-step and complex operations seamlessly.
-The agents can refine their responses based on feedback, learn from new information, and even interact with other AI agents to collaboratively solve complex tasks (Hong et al., 2024; Qian et al., 2023a; Xu et al., 2023; Lin et al., 2024).
+The agents can refine their responses based on feedback, learn from new information, and even interact with other AI agents to collaboratively solve complex tasks (Hong et al., 2024; Qian et al., 2023a; Xu et al., 2023; Lin et al., 2024b).
 Through prompting, agents can be assigned different roles (e.g., a developer or a tester), providing more domain-specific responses that help improve the answer (Hong et al., 2024; White et al., 2024; Shao et al., 2023).
-As their potential expands, LLM agents play a crucial role in advancing automation and boosting productivity in software development.
-In this paper, we explore using LLM agents to improve fault localization by emulating developers’ debugging process.
+Recent studies (Shinn et al., 2024; Renze and Guven, 2024; Pan et al., 2025) explore using the agent’s verbal reasoning result to guide iterative improvement (i.e., verbal reinforcement learning), which has shown promising improvement in downstream tasks.
+In verbal reinforcement learning, LLM agents receive natural language feedback, such as reasoning results or instructions, from other agents as a reward signal.
+This allows agents to learn and adapt their behavior based on human-like guidance, improving their learning process to solve complex tasks.
+As the capabilities of large language model (LLM) agents grow, they play an essential role in enhancing automation and increasing productivity in software development.
+They can assist in code generation (Nijkamp et al., 2022; Lin et al., 2024a; Gu, 2023), debugging (Lee et al., 2024; Kang et al., 2024), test case creation (Huang et al., 2024; Chen et al., 2024b), and automated refactoring (Pomian et al., 2024; Liu et al., 2025), enabling developers to streamline repetitive tasks and focus on higher-level design and problem-solving.
+Additionally, LLM agents can facilitate collaborative software engineering by acting as intelligent assistants in code reviews, documentation generation, and issue resolution, improving overall development efficiency (Lin et al., 2024a; He et al., 2024).
+In this paper, we explore using LLM agents to improve fault localization by emulating developers’ debugging process using verbal reinforcement learning.
 2.2. Related Work
@@
 Techniques like Grace (Lou et al., 2021) and GNET4FL (Qian et al., 2023b) utilize test coverage and source code structure for improved accuracy, while DepGraph (Rafi et al., 2024) refines these approaches by graph pruning and incorporating code change information, resulting in higher performance with reduced computational demands.
-Although these learning-based techniques show improved results, they require training data that may not be available to every project, and the training process can be expensive due to model complexity.
+Although these learning-based techniques show improved results, they require training data that may not be available to every project.
 LLM-Based Fault Localization.
@@
 LLMs have shown potential in identifying and fixing errors using program code and error logs (Achiam et al., 2023).
-However, one of the major challenges that LLMs face in fault localization is the token limitation issue.
-LLMs are restricted to a fixed number of tokens, ranging from 2,000 to 128,000 (OpenAI, 2024; Meta AI, 2024), which poses difficulties in handling large-scale software projects with long stack traces and extensive codebases.
-This limitation can lead to incomplete analyses, as critical context might be truncated or lost, forcing models to work with fragmented information, ultimately affecting the overall fault localization performance (Hadi et al., 2023; Hou et al., 2023).
-Moreover, LLM-based fault localization techniques often focus on localizing faults within small code snippets due to these context limitations.
-For example, LLMAO (Yang et al., 2024) uses lightweight bidirectional adapters on LLMs to generate suspiciousness scores for code lines, but only within a limited context of 128 lines of code.
-Similarly, Wu et al.
-(Wu et al., 2023) prompt ChatGPT with code and error logs to identify faulty lines, but these methods struggle to scale to larger software projects (Liu et al., 2024).
-AutoFL (Kang et al., 2024) is an LLM-based fault localization technique that detects faulty code locations and generates natural language explanations for bugs.
-However, it also faces difficulty handling large-scale projects due to LLM context length constraints and reduced performance on complex bugs requiring deeper repository exploration.
-To address these challenges, there is a growing need for approaches that enable LLMs to localize faults across entire software projects, ensuring they can handle larger inputs while maintaining accuracy and effectiveness.
-Thus, we propose LLM4FL, that leverages a divide-and-conquer approach to navigate through codebases and identify faults through multiple-agent collaboration.
+Some LLM-based fault localization techniques operate on small code snippets due to LLM’s token limitations.
+LLMAO (Yang et al., 2024) uses bidirectional adapters to score suspicious lines within a 128-line context.
+In contrast, Wu et al.
+(Wu et al., 2023) prompt ChatGPT with code and error logs but struggle to scale to large projects (Liu et al., 2024).
+Recent fault localization techniques employ different strategies to navigate large codebases and refine fault localization results (Kang et al., 2024; Zhang et al., 2024; Qin et al., 2024).
+Each technique consists of key components for retrieving failure-related classes or methods, navigating the code repository, and ranking suspicious methods.
+However, there are differences in how these techniques design the components.
+Table 1 provides a summary of the key differences between prior techniques and ours.
+AutoFL (Kang et al., 2024) retrieves class and method signatures related to the test failure, and prompts the LLM to decide which methods need future analysis (i.e., to examine the code).
+While this approach ensures the input size is less than the LLM’s token limitation, it does not consider method dependence in the analysis.
+Instead, it refines rankings by repeatedly running the fault localization process and using a majority voting mechanism to determine the most likely faulty methods.
+AgentFL/SoapFL (Qin et al., 2024) models fault localization as a structured operating procedure with controlled phases.
+It uses a document-guided approach, where the system creates (or refines) LLM-generated or developer-provided documentation for classes and methods, then navigates the codebase based on these documents.
+By iteratively scoring and ranking suspicious methods, it continuously narrows down the fault candidates.
+It integrates document-guided search and enhanced method documentation to navigate the codebase while using iterative LLM-based scoring and ranking to refine the set of suspicious methods.
+AutoCodeRover (Zhang et al., 2024) first extracts file/method names from an issue report.
+Then, based on the names, it searches for corresponding classes and methods.
+It is an iterative search process and may continue to search for the called methods in the repository based on LLM feedback.
+AutoCodeRover may also be integrated with SBFL to refine localization and guide patch generation.
+However, it does not explicitly rank faulty methods; instead, it selects a method from the retrieved context and generates a patch.
+In contrast to prior techniques, LLM4FL uses a structured, graph-based approach to analyze the code, enabling the LLM to systematically navigate the repository by considering caller-callee relationships.
+This eliminates LLM hallucination when searching and retrieving methods in the repository.
+Additionally, LLM4FL uses a divide-and-conquer strategy to mitigate token limitations by breaking down large methods into smaller, manageable groups, ensuring comprehensive analysis.
+Finally, LLM4FL refines fault localization results through verbal reinforcement learning, iteratively re-ranking suspicious methods for improved accuracy.
+We believe these contributions collectively enhance the accuracy and scalability of LLM-based fault localization.
 3. Methodology
-Figure 1. An overview of LLM4FL.
-Applying Large Language Models (LLMs) for fault localization presents specific challenges, such as token size limitations and performance issues with longer input contexts (Levy et al., 2024; Liu et al., 2024; Wu et al., 2023).
-Processing large datasets like test coverage and source code in a single query is impractical due to these constraints.
-Moreover, there are limited studies on effective prompting techniques for fault localization.
-To address these issues, we propose LLM4FL, an LLM-agent-based fault localization system.
-Figure 1 illustrates an overview of our technique.
-LLM4FL operates in four stages: 1) Analyzing the Covered Methods Using a Divide-and-Conquer Strategy, 2) Defining LLM Agent Roles and Tools, 3) Code Navigation through Agent Collaboration and Prompt Chaining, and 4) Self-Reflection and Chain-of-Thought Reasoning.
-First, the coverage data is divided into smaller, manageable groups to fit within the LLM’s token limits.
-Second, LLM agents are assigned specific roles and equipped with tools to analyze these groups.
-Third, LLM agents collaborate through prompt chaining to navigate the code autonomously and identify potential faults.
-Finally, the agents refine their analysis using self-reflection and chain-of-thought reasoning.
-Below, we provide more details on each of these stages.
-3.1. Analyzing the Covered Methods Using Divide-and-Conquer
-Prior studies in fault localization utilized dynamic execution data, such as coverage information from both failing and passing test cases, to identify faulty code locations (Lou et al., 2021; Li et al., 2019; Abreu et al., 2009; Rafi et al., 2024; Qian et al., 2023b).
-However, Large Language Models (LLMs) face challenges when processing large-scale coverage data due to inherent token size limitations (Kang et al., 2024).
-These limitations lead to performance degradation, extended processing times, and reduced accuracy when input tokens are truncated.
-To address these constraints, we propose a divide-and-conquer strategy, dividing the coverage data (i.e., containing the covered methods and the corresponding code) into manageable groups that the LLM can process within its token limits.
-We generate a list of covered methods using tools like GZoltar (Campos et al., 2012), focusing only on methods executed by failing tests.
-To manage the token limitation issue, we include only the specific method statements that were executed by the failing tests, excluding any uncovered portions of the code.
-Before doing a divide-and-conquer, we use a Spectrum-Based Fault Localization (SBFL) technique to sort the covered methods so that those more likely to be faulty can be grouped together.
-This approach is inspired by a recent study (Tyen et al., 2024) that demonstrates LLMs perform better when the order of instruction is carefully considered.
-Note that, regardless of the order, LLM4FL eventually analyzes every covered method.
-Specifically, we use Ochiai (Abreu et al., 2006) to sort the methods before the divide-and-conquer process.
-Ochiai is a widely used unsupervised spectrum-based fault localization technique because of its high efficiency and good fault localization accuracy (Abreu et al., 2006; Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021).
-Intuitively, Ochiai assigns a higher suspiciousness score to a statement that is executed more frequently by failing cases and less frequently by passing test cases.
-Ochiai is defined as:
-$Ochiai(a_{ef},a_{nf},a_{ep})=\frac{a_{ef}}{\sqrt{(a_{ef}+a_{nf})\times(a_{ef}+a_{ep})}},$
-where $a_{ef}$ is the number of failed test cases that execute a statement, $a_{nf}$ is the number of failed test cases that do not execute the statement, and $a_{ep}$ is the number of passed test cases that execute the statement.
-The result of Ochiai is a value between 0 and 1, with higher values indicating a higher likelihood that a particular statement is faulty.
-We then aggregate the scores for every statement in a method to get the method-level score.
-Once the covered methods are sorted, we divide the methods into smaller groups, ensuring that each group fits within the token limitation of the LLM.
-More formally, let $C$ represent the sorted covered methods, containing a sequence of pairs ($m$, $s$), where $m$ denotes a covered source method and $s$ denotes the corresponding code statements.
-The group length is determined using the token limitation specified in the official documentation of GPT-4o-mini (the LLM that we use in this study), which has an input context window of 128,000 tokens (OpenAI, 2024).
-The length is calculated as $TotalTokenLength/TokenLimitation$, ensuring each group remains within the model’s token limitation.
-Specifically, we divide $C$ into to sequence of $C_{1},C_{2},\ldots,C_{n}$ each $C_{i}$ contains the subset of pairs ($m_{i}$, $s_{i}$), and satisfies the constraint —$C_{i}$$<=TokenLimitation$.
-3.2. LLM Agents and the Available Tools
-LLM4FL emulates developers’ debugging process by implementing two LLM agents: 1) Tester and 2) Debugger. Below we discuss the details of these agents:
-Tester Agent.
-The tester agent’s main goal is to identify and prioritize suspicious methods.
-More formally, it navigates through each group $C_{i}$ of the covered methods.
-The Tester Agent’s input includes the sorted covered methods in each group $C_{i}$ and the corresponding test code, stack traces, and related test information from failing executions.
-To further reduce the input token size and allow LLM to focus on the relevant code, we only provide the portion of the test code leading to the failure as input.
-We analyze the stack trace to identify where the test failed and extract the code up to that point, omitting the rest.
-Figure 2 gives an example of the Tester Agent’s prompt.
-The Tester Agent is responsible for identifying suspicious methods likely related to faults in the system.
-To achieve this, it needs to 1) analyze methods executed by failing tests, 2) review relevant test code, and 3) examine stack trace data to pinpoint where failures occurred.
-To facilitate this process, we provide the Tester Agent with three tools.
-The tool get_covered_methods allows the agent to retrieve the source code of the methods executed during failing tests, get_test_code fetches the relevant test code, and get_stacktrace gathers stack trace data.
-Each tool has a unique name and description, which the agent uses to determine when to apply the tool during its analysis.
-The agent decides autonomously, based on its task, which tool to invoke by interpreting the tool’s description and relevance to the current stage of fault localization.
-Additionally, each method within the grouped data is assigned a unique ID during the dividing process, allowing the agent to reference and retrieve specific methods.
-After analyzing all groups, the Tester Agent generates a list of suspicious methods, denoted as $S_{i}\subseteq M_{i}$, where $M_{i}$ represents the methods in group $C_{i}$.
-This list of suspicious methods is then consolidated into a final list of method IDs for further analysis.
-Prompt Template for Initial Candidate Selection with Advanced Tools As an Intelligent Tester Agent, utilize the following tools to identify methods most likely related to the fault.
-Focus on analyzing the test code, stack trace, and covered methods to pinpoint suspicious or fault-inducing methods.
-Ensure the exact method_id is included in the final JSON output.
-Available Tools:
-get_test_code()       – Retrieves the complete test code.
-get_stacktrace()         – Access the stack trace of the failed test.
-get_covered_methods()   – Lists all methods executed during the test.
-Your Task: Use the insights from the tools and provide the potential suspicious methods in the following JSON format:
-        "method_ids": [1, 2, 3, ...]  // The potential suspicious method’s ids
-Figure 2. Enhanced Prompt for Initial Candidate Selection.
-Debugger Agent.
-Figure 3 gives an example of the Debugger Agent’s prompt.
-The Debugger Agent is responsible for thoroughly evaluating and ranking a given list of candidate methods $S$.
-To perform its analysis, the Debugger Agent relies on three tools designed to extract relevant information.
-get_test_code and get_stacktrace provide LLM’s requested information by retrieving the relevant test code and stack trace data, helping the Debugger Agent understand the failure scenarios associated with each method.
-Additionally, the get_method_body tool allows the agent to fetch the full source code of a method using its unique method ID, enabling the Debugger Agent to analyze each method’s source code.
-Once the Debugger Agent completes its analysis, it assigns a rank score $\sigma_{j}$ to each method $m_{j}$, representing the likelihood that the method is contributing to the fault.
-The candidate methods in $S$ are then ranked based on these scores, producing a ranked list $R$ that orders the methods from most to least suspicious:
-$R=\text{sort}(\{(m_{j},\sigma_{j}):m_{j}\in S\})$
-This final ranked list $R$ is output in JSON format for further evaluation and parsing.
-Prompt Template for Code Understanding and Fault Analysis As an Advanced Debugger Agent, use the following tools to analyze the provided method bodies based on the Tester Agent’s insights from the test and stack traces.
-The fault may propagate across multiple methods or be indirectly related to the original issue.
-Available Tools:
-get_test_code()       – Retrieve the entire test code for detailed analysis.
-get_stacktrace()         – Access the stack trace to trace method calls and identify potential faults.
-get_method_body()      – Retrieve the body of specific methods for further inspection.
-Task: Using the provided tools, analyze and focus on the following Methods: {suspicious methods} Rank these methods from most to least suspicious, and provide brief reasoning for each based on its behavior and potential involvement in the fault.
-IMPORTANT: Provide your output in the following JSON format:
-            "method_id": int,  // The most suspicious method’s id
-            "reasoning": string,  // Reasoning for its suspiciousness
-            "rank": int  // Rank of suspiciousness
-Figure 3. Enhanced Prompt for Code Understanding and Fault Analysis.
-3.3. Code Navigation Through Prompt Chaining
-Code navigation is a critical part of how developers trace faults in real-world scenarios.
-When debugging, developers often start by analyzing a specific method and then explore its caller or callee methods to better understand the overall logic and pinpoint where the fault might lie.
-Our approach emulates this process by implementing code navigation through prompt chaining, allowing the Tester and Debugger Agents to collaborate and progressively focus their analysis on the most relevant parts of the codebase.
-The process begins with the Tester Agent, which works through each group of the coverage data $C=\{C_{1},C_{2},\ldots,C_{n}\}$, where each group $C_{i}$ contains a subset of the covered methods $M_{i}$.
-The Tester Agent identifies suspicious methods within each group, producing a list $S_{i}\subseteq M_{i}$ of candidate methods.
-This step mimics how a developer might flag areas of interest in the code that seem related to the fault.
-After analyzing all the groups, the Tester Agent compiles the suspicious methods into a consolidated list:
-$S=\bigcup_{i=1}^{n}S_{i}$
-This list $S$ is passed to the Debugger Agent, initiating the prompt chaining process.
-The Debugger Agent takes the output from the Tester Agent and dives deeper into the suspicious methods.
-In this stage, the Debugger Agent navigates through each method in $S$, inspecting not only the method itself but also the related methods, such as caller and callee methods, if the agent believes further analysis can help identify faulty code location.
-The Debugger Agent emulates the manual process of tracing how faults propagate through different parts of the code.
-By leveraging its tools, the Debugger Agent retrieves information on how each method interacts with others and how it may contribute to the fault.
-If the Debugger Agent identifies that a method $m_{j+1}$ is called within $m_{j}$ and might require further examination, it can request the code for $m_{j+1}$ if it is part of the candidate list, ensuring deeper analysis where necessary.
-LLM4FL implements an LLM-based fault navigation system that allows the Debugger Agent to dynamically traverse the codebase and retrieve additional relevant methods for further inspection.
-Through this process, the Debugger Agent traces potential fault propagation pathways and evaluates how faults spread across methods.
-3.4. Self-Reflection and Chain-of-Thought for Improved Fault Localization Results
-LLM4FL uses two additional prompting techniques to further improve the fault localization results.
-Self-Reflection.
-In our approach, both the Tester and Debugger Agents engage in self-reflection to enhance the quality of their outputs.
-After completing their initial analysis, each agent operates in two phases: a generator phase and a reflector phase.
-During the generator phase, the agents produce their primary results, such as identifying suspicious methods or ranking them by suspiciousness.
-In the reflector phase, the agents critique their own work, offering feedback to refine their results and correct potential errors.
-While it has not yet been studied in the fault localization context, a recent study (Renze and Guven, 2024) found that self-reflection improves LLM’s performance in problem-solving tasks.
-Chain-of-Thought for Refining Fault Localization.
-We incorporate a chain-of-thought reasoning process to enhance fault localization.
-After generating a ranked list of suspicious methods, the LLMs propose potential fixes for the top-ranked methods.
-The agents then re-rank the methods by considering the generated fixes.
-This process allows the agents to think more deeply about the faults and reassess the ranking based on the insights gained from considering these fixes (Bao et al., 2024; Wang et al., 2024).
+Traditionally, developers rely on dynamic execution data to facilitate fault localization: (i) code coverage is used for narrowing down executed code that changes the behavior incorrectly, whereas (ii) stack trace provides developers with insight into the program execution when the exception gets thrown, helping them diagnose the issue more effectively.
+However, directly applying Large Language Models (LLMs) for fault localization is infeasible.
+This is due to the large size of software artifacts, which frequently exceed token limitations, leading to input truncation and degraded performance  (Levy et al., 2024; Liu et al., 2024; Wu et al., 2023).
+To mitigate this issue, we propose LLM4FL, an LLM-agent-based fault localization system, which adopts a divide-and-conquer algorithm, which first divides the large software artifacts, and employs multiple LLM-Agents to localize faults autonomously.
+Figure 1 provides an overview of our technique.
+LLM4FL consists of three LLM-Agents by using novel prompting techniques: (i) Context Extraction Agent, (ii) Debugger Agent, and (iii) Reviewer Agent to localize the fault iteratively.
+Specifically, the Context Extraction Agent utilizes a order-aware division and failure-reason guided prioritization prompting technique, which consists of two phases.
+In the division phase, a toolchain divides the large-scale code coverage data into small groups to fit within the LLM’s token limits.
+In the prioritization phase, the agent iteratively analyzes the divided code coverage, leveraging failed test cases and stack traces to identify potentially faulty methods.
+The Debugger Agent then performs graph-based retrieval-augmented code navigation prompting to locate code artifacts further to enhance the accuracy of fault ranking.
+Finally, the Reviewer Agent re-ranks the buggy method through verbal reinforcement learning prompting with the reviewer agent, ensuring more precise fault localization.
+Below, we provide more details on each of these phases.
+Figure 1.
+Overview of LLM4FL for Multi-Agent Fault Localization, illustrating how agents collaborate to analyze software artifacts, extract failure reasoning, perform graph-based retrieval-augmented code navigation, and rank faulty methods using verbal reinforcement learning.
+3.1. Context Extraction Agent
+This agent defines a order-aware division and failure-reason guided prioritization prompting technique to divide code-coverage and iteratively prioritize faulty methods using stack trace and test code within each division.
+We describe the technique below.
+3.1.1. Order-Aware Division Phase
+To mitigate token size limitation, our LLM4FL first performs a division phase, where it first runs the test case to extract the list of method-level code coverage using GZoltar (Campos et al., 2012), which is denoted as $C$, containing a sequence of pairs ($m$, $s$), where $m$ denotes the method and $s$ denotes the set of statements within method $m$.
+We then divide $C$ into $K$ group of sequences of $C_{1},C_{2},\ldots,C_{k}$, where $K=\left\lceil\frac{\text{Input Token Length}}{\text{Token Limitation}}\right\rceil$, and each subset satisfies $|C_{i}|\leq\lfloor\text{Token Limitation}\rfloor$ to ensure the data fits within the LLM’s context window.
+For example, given an Input Token Length of 500K tokens, we divide this by the Token Limitation for GPT-4o-mini, which has a limit of 128K tokens (OpenAI, 2024).
+This results in $K=4$, each with 128K token limit.
+Traditionally, the order in which divisions occur in divide-and-conquer algorithms does not impact the outcome.
+However, prior studies have shown that LLM may improve performance when the order of instructions is carefully considered  (Chen et al., 2024a).
+Inspired by this, we propose order-aware division, where we use a Spectrum-Based Fault Localization (SBFL) technique to sort $C$ based on their likelihood of being faulty.
+Specifically, we use Ochiai ranking to order $C_{1},C_{2},\ldots,C_{k}$, which is an efficient and unsupervised technique that assigns higher suspiciousness scores to statements that are executed more frequently by failing test cases and less frequently by passing ones (Abreu et al., 2006; Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021).
+3.1.2. Failure-Reason Guided Prioritization Phase
+To prioritize faulty methods from the divided code coverage, one approach is to provide the agent with the stack trace and test case.
+However, prior studies have shown that incorporating summarized description can improve accuracy (Stiennon et al., 2020; Roit et al., 2023).
+Thus, we propose failure-reason-guided fault prioritization, which follows two steps.
+First, the input tokens are summarized into a failure-reason representation, capturing the test purpose, expected output, and failure reason.
+Second, the divided code coverage is prioritized based on the reason for failure.
+Below, we discuss each step in detail.
+Figure 2.
+Context-Extraction Agent uses tool-chains to preprocess software artifacts for Lang-5 to emphasize the test failure context.
+For (i) stack-trace the agent prunes external libraries and (ii) test code, the agent prunes statements in the test code after the assertion failure.
+Generate failure-reason from code artifacts.
+Our failure-reason generation makes the following three observations: (i) Firstly, the stack traces are verbose and include calls to external libraries unrelated to the fault.
+(ii) Secondly, some statements within a test case are irrelevant to the failure, specifically statements after the first assertion failures.
+(iii) Finally, the test case may call other helper methods that trigger the test failure (Peng et al., 2022).
+Hence, as shown in Figure 2, to mitigate (i), the Context Extraction Agent uses PreprocessTrace tool-chain to prune external execution in the stack trace.
+To mitigate (ii) and (iii), Context Extraction Agent uses PreprocessTest tool-chain, which uses static analysis to build an interprocedural call graph to extract all helper test methods called by the test and prunes all the statements appearing in the test after the failure point.
+Given the preprocessed stack trace and test code, we prompt Context-Extraction Agents to generate failure-reason by answering test purpose, expected output, and failure reason.
+Below, we show a simplified example of the prompt:
+“Analyze the test body and stack trace to generate a test failure reason summarizing the issue with expected output, and potential faulty location.”
+## Input: ‘
+{Test Code, Stack Trace}
+## Response: ‘
+{Failure-Reason}
+An example of failure-reason generated for Lang-5 given stack trace and test code is:
+## Test Purpose:
+Test whether LocaleUtils.toLocale parses ’_GB’ as a locale string, expecting it to return a Locale object with an empty language and ’GB’ as the country.
+## Failure Reason:
+The actual output is an IllegalArgumentException due to an invalid locale format, as LocaleUtils.toLocale requires a language before the underscore.
+For example, ’_GB’ does not conform to the expected locale format.
+Update the test input to ’en_GB’ or modifying LocaleUtils.toLocale to handle ’_GB’ as a valid case.
+d
+Failure-reason guided prioritization.
+Given Failure-reason, the Context Extraction Agent analyzes the covered methods in each group, previously denoted as $\{C_{1},C_{2},\dots,C_{K}\}$, to prioritize methods that are most related to the failure.
+Formally, prioritizing each set $C_{i}$ results in a prioritized subset $C^{\prime}_{i}$, where: $C^{\prime}_{i}\subseteq C_{i},\forall i\in\{1,2,\dots,K\}$.
+The prioritized methods from all subsets are then combined into a final prioritized set: $C^{\prime}=\bigcup_{i=1}^{K}C^{\prime}_{i}$.
+This final prioritized list $C^{\prime}$ is just a union of the ordered subsets, meaning no additional processing is introduced.
+This union process is necessary as LLMs are highly sensitive to initial order (Chen et al., 2024a), yet analyzing entire code coverage poses a token limitation.
+By dividing coverage into groups, prioritizing each, and unionizing subsets, we extract important context that are necessary for the Debugger Agent for fault analysis and localization in the next step.
+3.2. Debugger Agent
+Previously, the Context Extraction Agent prioritized $C$ to subsets $C^{\prime}$ to mitigate the LLM’s token limitation when analyzing large coverage data.
+While this technique can adapt to any LLM and token input size, prioritization may miss methods in the coverage that are important for fault localization.
+Inspired by how developers analyze the call dependencies to understand program execution and identify faults, we propose a graph-based retrieval-augmented code navigation strategy (Graph-RAG).
+This approach enables the Debugger Agent to utilize call graphs and method bodies for effective repository navigation, implementation analysis, generation of failure reasoning, and ranking the methods based on their likelihood of causing errors.
+We will detail this technique below.
+3.2.1. Generating failure reasoning through call-graph-aware retrieval-augmented (Graph-RAG) code navigation.
+The agent first uses code-coverage $C$ to construct the inter-procedural call-graph database $G=(V,E)$, where $V$ is a set of methods in the call-graph, and $E$ is a set of edges representing the caller-callee relationship.
+This graph is used by the tool-chains, get_MethodBody and get_CallGraph to facilitate agent-driven navigation within the call graph.
+LLM4FL then uses the prompt template below to drive the agent:
+“Given (i) failure reasoning, (ii) stack trace, and (iii) test code, analyze, navigate, and enhance the failure reasoning for the given methods to identify faults.
+If additional implementation details are needed, extract the call graph and retrieve the caller or callee’s implementation.
+At the end, you should output the (i) all of the analyzed method, (ii) enhanced failure reasoning, (iii) fault ranking.”
+## Input:
+{Test Code, Stack Trace, Failure Reason, PrioritizedMethods}
+## Response:
+{Analyzed Method, Enhanced Failure Reason, Fault Ranking}
+The agent begins the analysis of methods selected sequentially from the prioritized $C^{\prime}$.
+For each method, the agent retrieves its body using get_MethodBody and analyzes its implementation.
+It then queries get_CallGraph to identify relevant callers or callees, deciding whether to retrieve their implementations to enhance its understanding of code behavior.
+Throughout the iterative process of retrieval and analysis, every method examined, regardless of whether it was prioritized or invoked by a prioritized method, is assigned a failure reason by the agent.
+This information is then saved in a set referred to as $R$.
+This failure reasoning is a natural language explanation of each method’s failure analyzed in the code navigation process.
+These results are stored as pairs in a set: $R=\{(m_{i},r_{i})\mid m_{i}\text{ is analyzed},\;i=1,\dots,n\}$.
+Here, $n$ denotes the number of methods for which failure reasoning was generated.
+Ranking methods based on failure reasoning results.
+Finally, given the failure reasoning $R$, the agent assigns each method $m_{i}\in\mathcal{R}$ an ordinal rank ($r^{*}_{i}$).
+Formally, the final ranked set $R^{*}$ is sorted in descending order based on the ordinal ranks:
+$R^{*}=\text{sort}\left(\{(m_{i},r_{i},r^{*}_{i})\mid m_{i}\in R\}\right)$
+where $r*_{i}$ denotes the ordinal rank that reflects each method’s likelihood of containing the fault, and $r_{i}$ explains the reasoning behind the method being faulty.
+The final ranked list $R^{*}$ is output in JSON format for subsequent analysis.
+3.3. Reviewer Agent
+Code review is important in software development to ensure software quality.
+Beyond traditional debugging, code review can help refine failure reasoning and improve fault localization.
+Inspired by this, we propose Reviewer Agent, which uses a novel prompting technique called Re-ranking through Verbal Reinforcement Learning to emulate a rigorous code-review process to refine fault localization.
+We describe the technique below.
+3.3.1. Re-ranking through Verbal Reinforcement Learning
+To refine the fault rankings $R^{*}$, Reviewer Agent adopts a verbal reinforcement learning process inspired by the Reflexion framework (Shinn et al., 2024).
+The agent iteratively critiques its rank trajectory by identifying inconsistencies, retrieving missing execution insights, and updating method prioritization to better reflect the failure’s root cause.
+After the iterations, it applies chain-of-thought to update the final ranking and return the fault localization results.
+“Given (i) failure reasoning, (ii) stack trace, and (iii) test code, analyze and critique the initial ranking of suspicious methods.
+Reflect on the reasoning quality and identify missing or superfluous details.
+If additional insights are needed, extract the call graph and retrieve the caller or callee’s implementation to refine the ranking.
+At the end, you should output (i) the revised set of analyzed methods, (ii) the improved failure reasoning, and (iii) the final fault ranking with a possible fix.”
+## Evaluating the Initial Ranking:
+{Test Code, Stack Trace, Ranked List with Reasonings}
+## Self-Critique and Refinement:
+- Identify missing insights and remove unnecessary information.
+- Retrieve method details using ‘get_MethodBody‘ if needed.
+- Extract the call graph with ‘get_CallGraph‘ to enhance fault reasoning.
+## Final Output:
+{Analyzed Methods, Enhanced Failure Reasoning, Fault Ranking}
+Evaluating and self-critiquing the initial ranking.
+The Reviewer Agent uses the initial ranking, $R^{*}$, provided by the Debugger Agent as the baseline trajectory for refinement.
+The agent enters a self-reflection phase, guided by Reflexion’s Self-Reflection Model ($M_{sr}$), to refine its ranking decisions through verbal reinforcement learning (Shinn et al., 2024).
+In this phase, the agent examines whether the initial ranking aligns with observed fault rankings by comparing the failure reasoning $r$, with caller-callee interactions and test coverage data, thereby identifying discrepancies such as misordered rankings or missing call dependencies.
+The agent generates natural language critiques and retrieves additional execution details if needed using tools such as get_MethodBody (to obtain complete method implementations) and get_CallGraph (to retrieve caller-callee relationships).
+The refined context is then incorporated in the Reviewer Agent’s iterative adjustment process.
+Ranking adjustment through trajectory optimization.
+After the initial self-critique phase, the Reviewer Agent refines its initial ranking by updating each method’s rank $r^{*}$ and revising its failure reasoning $r$.
+Conceptualized as a trajectory optimization problem within the Reflexion framework, the agent, acting as an Actor ($M_{a}$), leverages reinforcement cues from its Self-Reflection Model ($M_{sr}$) to systematically re-order methods based on the updated evidence of inter-method interactions and failure reasoning.
+Each iteration integrates new feedback to adjust the ranking until it stabilizes or a preset iteration limit is reached.
+Finalizing the ranking through chain-of-thought.
+After the iterative ranking adjustment, the Reviewer Agent finalizes the ranking one last time using chain-of-thought (Wang et al., 2024).
+To facilitate the thinking process, we ask the Reviewer Agent to generate a probable fix for every method in the final ranking by considering the updated ranking score $R^{*}$ and refined failure reasoning $R$.
+After generating all probable fixes, the agent then revisits all the information to do a final re-ranking.
+Formally, the final ranked list $R^{*}_{final}=\text{sort}\left(\{(m_{i},r^{*}_{i},r_{i},f_{i})\mid m_{i}\in C^{\prime}\}\right)$, where $f_{i}$ denotes the fix generated for method $m_{i}$.
+The final ranked list $R$, encoded in JSON format, provides a structured, machine-readable prioritization along with human-interpretable failure justifications.
 4. STUDY DESIGN AND RESULTS
-In this section, we first describe the study design and setup. Then, we present the motivation, approach, and results of the research questions.
+In this section, we first describe the study design and setup. Then, we present the answers to the research questions.
 Table 2.
@@
 To answer the RQs, we conducted the experiment on 675 faults across 14 projects from the Defects4J benchmark (V2.0.0) (Just et al., 2014).
-Defects4J provides a controlled environment to reproduce faults collected from projects of various types and sizes.
+Defects4J provides a controlled environment to reproduce faults collected from various types and sizes of projects.
 Defects4J is widely used in prior automated fault localization research (Lou et al., 2021; Sohn and Yoo, 2017; Chen et al., 2022; Zhang et al., 2017).
@@
 Table 2 gives detailed information on the projects and faults we use in our study.
-In total, the faults have over 1.4K fault-triggering tests (i.e., failing tests that cover the fault).
+The faults have over 1.4K fault-triggering tests (i.e., failing tests that cover the fault).
 The sizes of the studied projects range from 2K to 90K lines of code.
@@
 To collect test coverage data and compute results for baseline techniques, we utilized Gzoltar (Campos et al., 2012), an automated tool that executes tests and gathers coverage information.
-For the LLM-based components, we employed OpenAI’s GPT-4o mini, a more cost-effective and capable alternative to GPT-3.5 Turbo (OpenAI, 2024).
-LangChain v0.2 was used to streamline the development of LLM4FL, facilitating the integration of language models with external tools and enhancing the system’s overall functionality (Langchain, 2024a).
-To implement the self-reflection technique, we leveraged the LangGraph framework, which enabled graph-based reasoning and decision-making processes (Langchain, 2024b).
-To minimize the variations in the output, we set the temperature parameter to 0 during model inference.
-4.1. RQ1: How does LLM4FL perform compared with other fault localization techniques?
+For the LLM-based components, we employed OpenAI’s gpt-4o-mini-2024-07-18, a more cost-effective yet capable LLM (OpenAI, 2024).
+We used LangChain to develop LLM4FL (Langchain, 2024).
+We designed the prompts to be concise to minimize token usage and to allow more room for analysis-related information and code.
+To reduce the variations in the output, we set the temperature parameter to 0 during model inference.
+RQ1: How does LLM4FL perform compared with other fault localization techniques?
 Motivation. In this RQ we evaluate the fault localization accuracy of LLM4FL by comparing with various baseline techniques.
 Approach.
-We compare LLM4FL’s fault localization accuracy with five baselines: Ochiai (Abreu et al., 2006), DeepFL (Li et al., 2019), Grace (Lou et al., 2021), DepGraph (Rafi et al., 2024), and AutoFL (Kang et al., 2024).
-Ochiai (Abreu et al., 2006) is widely recognized in fault localization research for its high efficiency and accuracy, making it a common baseline for comparison (Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021; Rafi et al., 2024).
+We compare LLM4FL’s fault localization accuracy with five baselines representing different methodological families: Ochiai (statistical) (Abreu et al., 2006), DeepFL (deep neural network) (Li et al., 2019), Grace (graph neural network) (Lou et al., 2021), DepGraph (graph neural network) (Rafi et al., 2024), AutoFL (LLM) (Kang et al., 2024), and AgentFL/SoapFL (Qin et al., 2024) (LLM).
+Ochiai (Abreu et al., 2006) is a widely recognized statistical fault localization technique known for its high efficiency, making it a common baseline for comparison (Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021; Rafi et al., 2024).
 As such, we use Ochiai to rank the methods during the segmentation process and include it as a baseline for accuracy comparison.
@@
 We follow the study (Li et al., 2019) to implement DeepFL and include the SBFL scores from 34 techniques, code complexity, and textual similarities as part of the features for the deep learning model.
-Grace (Lou et al., 2021) is one of the first fault localization techniques based on graph neural networks (GNN) that represents code as a graph and uses a gated graph neural network to rank the faulty methods.
-Since then, GNN-based techniques have shown superior fault localization accuracy compared to traditional techniques.
-DepGraph (Rafi et al., 2024) is a GNN-based technique that further improves Grace by enhancing code representation in a graph using interprocedural call graph analysis for graph pruning and integrating historical code change information in the graph.
-AutoFL is a LLM-based fault localization technique.
-It begins by providing the LLM with information about a failing test and descriptions of specific methods that can be used to navigate the codebase.
-The LLM then interacts with these methods to gather relevant information, such as covered classes, methods, and code snippets.
-Scoring and ranking candidate methods (depicted as black rectangles) based on five AutoFL prediction outcomes.
-The methods are ranked by assigning scores based on multiple runs of AutoFL.
-In each run, a method’s score is the inverse of the total number of predicted methods, and these scores are averaged across all runs.
-Methods are then ranked in descending order of their average scores, with earlier predictions used to break any ties.
-While the original paper used OpenAI’s GPT-3.5-turbo-0613 model for their experiments, for our evaluation, we are using the latest lightweight GPT-4o mini model to perform AutoFL’s experiments.
-Table 2. Fault localization accuracy in terms of Top-1, 3, 5, and 10. The numbers in the parenthesis show the number of faults in each project. The result continues in Table 3.
-        Project (# faults)         Techniques         Top-1         Top-3         Top-5         Top-10
-        Cli (39)         Ochiai         3.0         5.0         10.0         18.0
-        DeepFL         11.0         21.0         24.0         28.0
-        Grace         14.0         24.0         26.0         30.0
-        DepGraph         17.0         24.0         27.0         34.0
-        AutoFL         12.0         19.0         19.0         20.0
-        LLM4FL         16.0         21.0         23.0         24.0
-        Closure (174)         Ochiai         20.0         39.0         70.0         72.0
-        DeepFL         46.0         61.0         92.0         99.0
-        Grace         51.0         78.0         102.0         121.0
-        DepGraph         60.0         99.0         126.0         148.0
-        AutoFL         45.0         58.0         65.0         82.0
-        LLM4FL         52.0         77.0         102.0         118.0
-        Codec (18)         Ochiai         3.0         12.0         17.0         17.0
-        DeepFL         5.0         10.0         12.0         16.0
-        Grace         6.0         11.0         13.0         17.0
-        DepGraph         7.0         10.0         14.0         16.0
-        AutoFL         12.0         14.0         14.0         16.0
-        LLM4FL         9.0         13.0         13.0         13.0
-        Collections (4)         Ochiai         1.0         1.0         2.0         2.0
-        DeepFL         1.0         1.0         2.0         2.0
-        Grace         1.0         1.0         2.0         2.0
-        DepGraph         1.0         2.0         2.0         2.0
-        AutoFL         1.0         1.0         1.0         1.0
-        LLM4FL         1.0         1.0         1.0         1.0
-        Compress (47)         Ochiai         5.0         12.0         17.0         29.0
-        DeepFL         22.0         27.0         31.0         38.0
-        Grace         23.0         29.0         34.0         42.0
-        DepGraph         25.0         33.0         36.0         45.0
-        AutoFL         23.0         33.0         34.0         35.0
-        LLM4FL         23.0         32.0         34.0         34.0
-        Csv (16)         Ochiai         3.0         8.0         10.0         12.0
-        DeepFL         7.0         8.0         9.0         11.0
-        Grace         6.0         8.0         10.0         12.0
-        DepGraph         8.0         9.0         12.0         13.0
-        AutoFL         5.0         11.0         12.0         14.0
-        LLM4FL         8.0         10.0         10.0         10.0
-        Gson (16)         Ochiai         4.0         9.0         9.0         12.0
-        DeepFL         8.0         11.0         12.0         12.0
-        Grace         11.0         13.0         14.0         15.0
-        DepGraph         14.0         15.0         16.0         16.0
-        AutoFL         5.0         7.0         10.0         10.0
-        LLM4FL         11.0         14.0         14.0         14.0
-Table 3. Continued from Table 2. Fault localization accuracy in terms of Top-1, 3, 5, and 10. The numbers in the parenthesis show the number of faults in each project.
-        Project (# faults)         Techniques         Top-1         Top-3         Top-5         Top-10
-        JacksonCore (26)         Ochiai         6.0         11.0         13.0         14.0
-        DeepFL         5.0         5.0         9.0         10.0
-        Grace         9.0         13.0         14.0         15.0
-        DepGraph         12.0         15.0         15.0         16.0
-        AutoFL         10.0         17.0         17.0         17.0
-        LLM4FL         12.0         13.0         15.0         15.0
-        JacksonXml (6)         Ochiai         0.0         0.0         0.0         0.0
-        DeepFL         3.0         3.0         4.0         5.0
-        Grace         3.0         3.0         4.0         5.0
-        DepGraph         4.0         5.0         5.0         5.0
-        AutoFL         2.0         2.0         2.0         3.0
-        LLM4FL         4.0         4.0         4.0         4.0
-        Jsoup (93)         Ochiai         15.0         40.0         48.0         57.0
-        DeepFL         33.0         39.0         46.0         49.0
-        Grace         40.0         64.0         72.0         77.0
-        DepGraph         53.0         73.0         78.0         83.0
-        AutoFL         36.0         52.0         52.0         54.0
-        LLM4FL         41.0         56.0         60.0         60.0
-        Lang (64)         Ochiai         25.0         45.0         51.0         59.0
-        DeepFL         42.0         53.0         55.0         57.0
-        Grace         43.0         53.0         57.0         58.0
-        DepGraph         48.0         55.0         60.0         61.0
-        AutoFL         40.0         57.0         60.0         60.0
-        LLM4FL         48.0         55.0         58.0         58.0
-        Math (106)         Ochiai         23.0         52.0         62.0         82.0
-        DeepFL         52.0         81.0         90.0         95.0
-        Grace         64.0         79.0         92.0         97.0
-        DepGraph         72.0         92.0         97.0         102.0
-        AutoFL         53.0         81.0         87.0         94.0
-        LLM4FL         68.0         87.0         92.0         94.0
-        Time (26)         Ochiai         6.0         12.0         13.0         16.0
-        DeepFL         12.0         15.0         18.0         20.0
-        Grace         11.0         16.0         20.0         21.0
-        DepGraph         17.0         20.0         21.0         22.0
-        AutoFL         13.0         18.0         21.0         22.0
-        LLM4FL         14.0         21.0         22.0         23.0
-        Mockito (38)         Ochiai         7.0         14.0         18.0         23.0
-        DeepFL         10.0         18.0         23.0         26.0
-        Grace         16.0         24.0         26.0         29.0
-        DepGraph         21.0         29.0         32.0         34.0
-        AutoFL         18.0         23.0         29.0         29.0
-        LLM4FL         21.0         22.0         26.0         27.0
-        Total (675)         Ochiai         121.0         260.0         340.0         413.0
-        DeepFL         257.0         353.0         427.0         468.0
-        Grace         298.0         416.0         486.0         541.0
-        DepGraph         359.0         481.0         541.0         597.0
-        AutoFL         275.0         393.0         423.0         457.0
-        LLM4FL         328.0         426.0         474.0         495.0
+Grace (Lou et al., 2021) utilized graph neural networks (GNN).
+It represents code as a graph and uses a gated graph neural network to rank the faulty methods.
+DepGraph (Rafi et al., 2024) is another GNN-based technique that further improves Grace by enhancing code representation in a graph using interprocedural call graph analysis for graph pruning and integrating historical code change information.
+AutoFL (Kang et al., 2024) is an LLM-based fault localization approach that provides the LLM with a failing test and method descriptions to gather relevant coverage data, then repeats the process to assign inverse scores and rank candidate methods by averaging results across runs.
+SoapFL (also known as AgentFL(Qin et al., 2024) organizes fault localization into structured phases, including test failure comprehension, code navigation, and iterative LLM-based scoring, leveraging enhanced method documentation to refine suspicious methods.
+While both AutoFL and SoapFL used OpenAI’s GPT-3.5 for their experiments, for our evaluation, we adapted both techniques to use the same LLM version as LLM4FL (i.e., gpt-4o-mini-2024-07-18) for comparison.
+Table 3. Fault localization accuracy (Top-1, 3, 5, and 10) for 675 faults from Defect4J V2.0.0. The numbers in the parenthesis show the percentage difference compared to LLM4FL.
+Techniques Top-1 Top-3 Top-5 Top-10
+Ochiai 121 (169.42%) 260 (63.46%) 340 (39.41%) 413 (20.10%)
+DeepFL 257 (26.85%) 353 (20.40%) 427 (11.01%) 468 (5.98%)
+Grace 298 (9.40%) 416 (2.16%) 486 (-2.47%) 541 (-8.32%)
+DepGraph 359 (-9.19%) 481 (-11.64%) 541 (-12.38%) 597 (-16.92%)
+AutoFL 275 (18.55%) 393 (8.14%) 423 (12.06%) 457 (8.53%)
+SoapFL/AgentFL 311 (4.82%) 414 (2.66%) 455 (4.18%) 478 (3.77%)
+LLM4FL 326 425 474 496
 Results.
-LLM4FL outperforms the LLM-based baseline, AutoFL, by achieving a much higher Pass@1 (328 v.s.
-275).
-Table 2 and 3 show the fault localization results of LLM4FL and the baseline techniques.
-Between the two LLM-based techniques, LLM4FL achieves a better Top@N than AutoFL when N is 1, 3, 5, and 10.
-In the Top-1 metric, LLM4FL locates the correct fault in 328 cases, compared to AutoFL’s 275, representing a 19.27% improvement.
-Similarly, in the Top-3 and Top-5 metrics, LLM4FL achieves scores of 426 and 474, respectively, outperforming AutoFL’s results of 393 and 423.
-Even in the broader Top-10 metric, LLM4FL shows an 8.32% advantage, with a score of 495 compared to AutoFL’s 457.
-These numbers highlight LLM4FL’s enhanced ability to pinpoint faulty methods more accurately and efficiently, reinforcing its superiority over AutoFL for fault localization tasks.
-LLM4FL shows higher Top-1 and Top-3 compared to all other non-LLM-based techniques, except for DepGraph.
-For the Top-1 metric, LLM4FL scores 328, which is 171.07% higher than Ochai’s score of 121, 27.63% higher than DeepFL’s score of 257, and 10.07% better than Grace’s result of 298.
+LLM4FL outperforms the LLM-based baselines, AutoFL and AgentFL, by achieving a 185.55% and 4.82% improvement in Pass@1, respectively.
+Tables 3 show the fault localization results of LLM4FL and the baseline techniques.
+Among the three LLM-based techniques, LLM4FL achieves a better Top@N across all values of N.
+In the Top-1 metric, LLM4FL locates the correct fault in 326 cases, compared to AutoFL’s 275 and AgentFL’s 311, representing an 18.55% and 4.82% improvement, respectively.
+Similarly, in Top-3, Top-5, and Top-10, LLM4FL an improvement between 8.14% to 12.06% over AutoFL and 2.66% to 4.18% for AgentFL.
+These numbers highlight LLM4FL’s ability to pinpoint faulty methods more accurately.
+For cost, we compute the total dollars spent by multiplying each approach’s token usage by the per-million-token price (e.g., \$0.150 for input, \$0.600 for output).
+SOAPFL averages about \$0.055 per bug, AutoFL averages about \$0.065, while LLM4FL is at around \$0.050 per bug making our approach comparably cost-effective.
+LLM4FL shows higher Top-1 and Top-3 compared to most other non-LLM-based techniques.
+For the Top-1 metric, LLM4FL scores 326, which is 169.42% higher than Ochia’s score of 121, 26.84% higher than DeepFL’s score of 257, and 9.39% better than Grace’s result of 298.
 One exception is DepGraph, which achieves a Top-1 of 359, 8.64% higher than LLM4FL.
-As the range expands to Top-3 and beyond, LLM4FL continues to demonstrate its robustness, significantly outperforming DeepFL and maintaining strong performance alongside Grace.
-It is important to note that DeepFL, Grace, and DepGraph are supervised methods trained and evaluated using leave-one-out cross-validation, following the original study protocols.
-Despite LLM4FL being a zero-shot approach without task-specific training, its ability to perform competitively with these supervised techniques further highlights its effectiveness.
-The fact that LLM4FL achieves strong results without the need for extensive training data or cross-validation underscores its potential and sheds light on future research directions, especially when training models on specific data sets may not always be feasible.
-LLM4FL outperforms the other LLM-based technique, AutoFL, by achieving a 19.27% higher Top-1 and delivering competitive results compared to supervised techniques like DeepFL and Grace, even without task-specific training.
-The findings highlight the potential of using LLM4FL for zero-shot fault localization.
-4.2. RQ2: How do different components affect LLM4FL’s fault localization accuracy?
+As the range expands to Top-3 and beyond, LLM4FL demonstrates its robustness, significantly outperforming DeepFL and maintaining strong performance alongside Grace.
+LLM-based techniques have several advantages over traditional techniques such as DepGraph.
+First, techniques like LLM4FL leverage pre-trained LLMs in a zero-shot setting, which can be easily applied to systems with insufficient training data.
+Second, techniques like DepGraph require additional model training that can take days.
+LLM-based techniques leverage generic pre-trained LLMs without specific fine-tuning or re-training.
+Finally, LLM-based techniques can explain the decision, which helps with adaption (Kang et al., 2024; Qin et al., 2024).
+Hence, LLM-based techniques have a strong potential to enhance fault localization by offering greater adaptability, reducing training overhead, and providing interpretable explanations.
+LLM4FL achieves a higher Top-1 compared to other LLM-based techniques, AutoFL and AgentFL, by 18.55% and 4.82%, respectively.
+It also has competitive results compared to supervised techniques like DeepFL and Grace by leveraging a pre-trained LLM in a zero-shot setting.
+RQ2: How do different components in LLM4FL affect the fault localization accuracy?
 Motivation.
-In this research question, we study the influence of each individual component within LLM4FL on its fault localization performance.
-LLM4FL employs a combination of advanced techniques, including coverage splitting to handle token limitations, prompt chaining for fault navigation, self-reflection for iterative refinement of the agents’ decisions, and chain-of-thought reasoning for re-ranking suspicious methods.
-Each of these components plays a distinct role in the overall process.
-To study their effects, we investigate how the removal of any one of these components affects the LLM4FL’s ability to accurately localize faults.
-By isolating the effects of each technique, we can better understand which components are critical to maintaining or improving the overall accuracy of the LLM4FL and provide insights to improving future LLM-based fault localization techniques.
-Approach.
-To evaluate the impact of each component, we designed four experimental configurations: LLM4FLw/o CodeNavigation, LLM4FLw/o Divide&Conquer, LLM4FLw/o Reflection, and LLM4FLw/o CoT.
-For each configuration, one component is systematically removed, and we compare the result with the full-featured baseline of LLM4FL.
-This allows us to directly measure how much each component contributes to the fault localization process.
-LLM4FLw/o CodeNavigation removes the prompt chaining mechanism, which essentially means the agents no longer collaborate do fault navigation through multiple rounds of reasoning.
-Instead, the LLM4FL uses a single prompt to perform fault localization without iterative agent communication and fault navigation.
-This configuration tests whether the multi-step agent collaboration improves the ranking and selection of faulty methods or if a single round of communication is sufficient.
-LLM4FLw/o Divide&Conquer removes the grouping of the covered methods, giving the agents the entire coverage data at once instead of dividing it into smaller, manageable groups.
+LLM4FL employs a number of components, and each of these components plays a distinct role in the overall process.
+Hence, in this RQ, we conduct an ablation study by removing each component separately and study their impact on fault localization accuracy.
+The findings may inspire future studies on adapting the components for similar tasks.
+Approach. To evaluate the impact of each component, we designed four different configurations:
+LLM4FLw/o CodeNav removes the code navigation mechanism, which essentially means the Debugger agent no longer does fault navigation by retrieving the call graphs.
+Instead, the LLM4FL uses a single prompt to perform fault localization without fault navigation.
+This configuration tests whether using the caller-callee information improves the ranking and selection of faulty methods or if a single round of analysis is sufficient.
+LLM4FLw/o Division removes the order-aware division of the covered methods, giving the agents the entire coverage data at once instead of dividing it into smaller, manageable groups.
 Coverage segmentation addresses token limitations in LLMs, so removing it explores the impact of feeding the full dataset to the agents in one step.
 We aim to see how handling large amounts of data in a single input influences the fault localization result, as it may overwhelm the model or reduce precision.
-LLM4FLw/o Reflection removes the self-reflection technique, which is used to allow agents to review and refine their initial candidate selection and ranking.
-Without this self-reflection step, the agents rely solely on their initial assessments without iterative improvements.
-LLM4FLw/o CoT disables the chain-of-thought reasoning process, which asks the agents to think deeper by asking LLMs to generate potential fixes and re-rank the suspicious methods.
-The chain-of-thought approach is designed to enhance logical reasoning and ensure that intermediate results are critically evaluated and refined.
-Table 4. Impacts of removing different components in LLM4FL on Top-1, 3, 5, and 10. The numbers in the parenthesis show the percentage changes compared to LLM4FL with all the components.
-       Project (# faults)        Techniques        Top-1        Top-3        Top-5        Top-10
-       Cli (39)        LLM4FL        16        21        23        24
-       LLM4FLw/o CodeNavigation        11 (-31.25%)        19 (-9.52%)        21 (-8.7%)        21 (-12.5%)
-       LLM4FLw/o Divide&Conquer        11 (-31.25%)        17 (-19.05%)        17 (-26.09%)        19 (-20.83%)
-       LLM4FLw/o Reflection        11 (-31.25%)        19 (-9.52%)        21 (-8.7%)        21 (-12.5%)
-       LLM4FLw/o CoT        16 (0.0%)        21 (0.0%)        23 (0.0%)        24 (0.0%)
-       Closure (174)        LLM4FL        52        77        102        118
-       LLM4FLw/o CodeNavigation        32 (-38.46%)        45 (-41.56%)        53 (-48.04%)        53 (-55.08%)
-       LLM4FLw/o Divide&Conquer        27 (-48.08%)        40 (-48.05%)        49 (-51.96%)        50 (-57.63%)
-       LLM4FLw/o Reflection        50 (-3.85%)        74 (-3.9%)        92 (-9.8%)        109 (-7.63%)
-       LLM4FLw/o CoT        52 (0.0%)        77 (0.0%)        102 (0.0%)        118 (0.0%)
-       Codec (18)        LLM4FL        9        13        13        13
-       LLM4FLw/o CodeNavigation        8 (-11.11%)        12 (-7.69%)        13 (0.0%)        13 (0.0%)
-       LLM4FLw/o Divide&Conquer        8 (-11.11%)        11 (-15.38%)        12 (-7.69%)        13 (0.0%)
-       LLM4FLw/o Reflection        9 (0.0%)        12 (-7.69%)        13 (0.0%)        13 (0.0%)
-       LLM4FLw/o CoT        9 (0.0%)        13 (0.0%)        13 (0.0%)        13 (0.0%)
-       Compress (47)        LLM4FL        23        32        34        34
-       LLM4FLw/o CodeNavigation        23 (0.0%)        32 (0.0%)        34 (0.0%)        34 (0.0%)
-       LLM4FLw/o Divide&Conquer        23 (0.0%)        28 (-12.5%)        30 (-11.76%)        31 (-8.82%)
-       LLM4FLw/o Reflection        23 (0.0%)        32 (0.0%)        35 (2.94%)        35 (2.94%)
-       LLM4FLw/o CoT        23 (0.0%)        32 (0.0%)        34 (0.0%)        34 (0.0%)
-       Csv (16)        LLM4FL        8        10        10        10
-       LLM4FLw/o CodeNavigation        8 (0.0%)        8 (-20.0%)        9 (-10.0%)        9 (-10.0%)
-       LLM4FLw/o Divide&Conquer        8 (0.0%)        8 (-20.0%)        9 (-10.0%)        9 (-10.0%)
-       LLM4FLw/o Reflection        7 (-12.5%)        10 (0.0%)        10 (0.0%)        10 (0.0%)
-       LLM4FLw/o CoT        8 (0.0%)        9 (-10.0%)        10 (0.0%)        10 (0.0%)
-       Gson (16)        LLM4FL        11        14        14        14
-       LLM4FLw/o CodeNavigation        9 (-18.18%)        15 (7.14%)        15 (7.14%)        15 (7.14%)
-       LLM4FLw/o Divide&Conquer        10 (-9.09%)        14 (0.0%)        15 (7.14%)        15 (7.14%)
-       LLM4FLw/o Reflection        11 (0.0%)        12 (-14.29%)        13 (-7.14%)        14 (0.0%)
-       LLM4FLw/o CoT        11 (0.0%)        14 (0.0%)        14 (0.0%)        14 (0.0%)
-       JacksonCore (26)        LLM4FL        12        13        15        15
-       LLM4FLw/o CodeNavigation        9 (-25.0%)        13 (0.0%)        14 (-6.67%)        14 (-6.67%)
-       LLM4FLw/o Divide&Conquer        9 (-25.0%)        13 (0.0%)        13 (-13.33%)        13 (-13.33%)
-       LLM4FLw/o Reflection        10 (-16.67%)        13 (0.0%)        13 (-13.33%)        13 (-13.33%)
-       LLM4FLw/o CoT        12 (0.0%)        13 (0.0%)        15 (0.0%)        15 (0.0%)
-       JacksonXml (6)        LLM4FL        4        4        4        4
-       LLM4FLw/o CodeNavigation        2 (-50.0%)        3 (-25.0%)        3 (-25.0%)        3 (-25.0%)
-       LLM4FLw/o Divide&Conquer        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)
-       LLM4FLw/o Reflection        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)        2 (-50.0%)
-       LLM4FLw/o CoT        3 (-25.0%)        4 (0.0%)        4 (0.0%)        4 (0.0%)
-       Jsoup (93)        LLM4FL        41        56        60        60
-       LLM4FLw/o CodeNavigation        40 (-2.44%)        55 (-1.79%)        58 (-3.33%)        58 (-3.33%)
-       LLM4FLw/o Divide&Conquer        36 (-12.2%)        50 (-10.71%)        51 (-15.0%)        51 (-15.0%)
-       LLM4FLw/o Reflection        38 (-7.32%)        53 (-5.36%)        54 (-10.0%)        54 (-10.0%)
-       LLM4FLw/o CoT        40 (-2.44%)        55 (-1.79%)        59 (-1.67%)        60 (0.0%)
-       Lang (64)        LLM4FL        48        55        58        58
-       LLM4FLw/o CodeNavigation        42 (-12.5%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
-       LLM4FLw/o Divide&Conquer        43 (-10.42%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
-       LLM4FLw/o Reflection        42 (-12.5%)        54 (-1.82%)        59 (1.72%)        59 (1.72%)
-       LLM4FLw/o CoT        48 (0.0%)        55 (0.0%)        59 (1.72%)        59 (1.72%)
-       Math (106)        LLM4FL        68        87        92        94
-       LLM4FLw/o CodeNavigation        59 (-13.24%)        79 (-9.2%)        85 (-7.61%)        85 (-9.57%)
-       LLM4FLw/o Divide&Conquer        49 (-27.94%)        68 (-21.84%)        70 (-23.91%)        78 (-17.02%)
-       LLM4FLw/o Reflection        57 (-16.18%)        80 (-8.05%)        85 (-7.61%)        85 (-9.57%)
-       LLM4FLw/o CoT        68 (0.0%)        87 (0.0%)        91 (-1.09%)        94 (0.0%)
-       Mockito (38)        LLM4FL        21        22        26        27
-       LLM4FLw/o CodeNavigation        18 (-14.29%)        23 (4.55%)        23 (-11.54%)        23 (-14.81%)
-       LLM4FLw/o Divide&Conquer        15 (-28.57%)        20 (-9.09%)        21 (-19.23%)        21 (-22.22%)
-       LLM4FLw/o Reflection        18 (-14.29%)        20 (-9.09%)        20 (-23.08%)        22 (-18.52%)
-       LLM4FLw/o CoT        21 (0.0%)        22 (0.0%)        26 (0.0%)        27 (0.0%)
-       Time (26)        LLM4FL        14        21        22        23
-       LLM4FLw/o CodeNavigation        12 (-14.29%)        20 (-4.76%)        22 (0.0%)        22 (-4.35%)
-       LLM4FLw/o Divide&Conquer        10 (-28.57%)        16 (-23.81%)        17 (-22.73%)        20 (-13.04%)
-       LLM4FLw/o Reflection        12 (-14.29%)        19 (-9.52%)        19 (-13.64%)        22 (-4.35%)
-       LLM4FLw/o CoT        14 (0.0%)        22 (4.76%)        23 (4.55%)        23 (0.0%)
-       Total (675)        LLM4FL        327        425        473        494
-       LLM4FLw/o CodeNavigation        273 (-16.51%)        378 (-11.06%)        409 (-13.53%)        409 (-17.21%)
-       LLM4FLw/o Divide&Conquer        251 (-23.24%)        341 (-19.76%)        365 (-22.83%)        381 (-22.87%)
-       LLM4FLw/o Reflection        290 (-11.31%)        400 (-5.88%)        436 (-7.82%)        459 (-7.09%)
-       LLM4FLw/o CoT        325 (-0.61%)        424 (-0.24%)        473 (0.0%)        495 (0.2%)
+LLM4FLw/o Reflexion removes the verbal reinforcement learning technique, which is used to allow agents to review and refine their initial ranking.
+Without this step, the agents rely solely on their initial assessments without iterative improvements.
+Table 4. Impacts of removing different components in LLM4FL on Top-1, 3, 5, and 10. The numbers in the parentheses show the percentage changes compared to LLM4FL with all the components.
+Techniques Top-1 Top-3 Top-5 Top-10
+LLM4FL 327 425 473 494
+LLM4FLw/o CodeNav 273 (-16.51%) 378 (-11.06%) 409 (-13.53%) 409 (-17.21%)
+LLM4FLw/o Division 251 (-23.24%) 341 (-19.76%) 365 (-22.83%) 381 (-22.87%)
+LLM4FLw/o Reflexion 290 (-11.31%) 400 (-5.88%) 436 (-7.82%) 459 (-7.09%)
 Results.
-While all components help improve the results, including coverage grouping and prompt chaining provide the largest improvement to fault localization results (23% and 17% in Top-1).
+While all components help improve the results, including coverage division and code navigation provide the largest improvement to fault localization results (23% and 17% in Top-1).
 Table 4 shows the Top-1, 3, 5, and 10 scores when each component is removed.
-Removing coverage splitting has the largest overall impact across all scores, reducing Top-N by 19% to 23%.
+Removing coverage divisions has the largest overall impact across all scores, reducing Top-N by 19% to 23%.
 Removing prompt chaining has the second largest impact (11% to 17%).
@@
 Future research should consider these techniques when designing fault localization techniques.
-Although there is no oracle during the fault localization process, asking LLMs to self-reflection still helps improve the overall Top-1 by 11%.
+Although there is no oracle during the fault localization process, asking LLMs to self-reflect still helps improve the overall Top-1 by 11%.
 LLMs often suffer from hallucinations, especially when there is a lack of feedback from external oracles (Xu et al., 2024; Huang et al., 2023).
-Even though we did not provide any groundtruth or external feedback to LLM, we find that self-reflection is still effective in improving fault localization results.
-Self-reflect brings 6% to 11% improvement across the Top-N metrics.
+Even though we did not provide any groundtruth or external feedback to LLM, we found that reflexion is still effective in improving fault localization results.
+We speculate that verbal reinforcement learning helps the model improve its results by revisiting the suspicious methods it ranked earlier, creating a feedback loop.
+In this process, we analyze the results for each group of methods and combine these group-specific results into a final ranking.
+The model can then spot mistakes or gaps in its logic, leading to better results.
+Self-reflection brings 6% to 11% improvement across the Top-N metrics.
 Our finding suggests that future studies should consider self-reflection even if there is no external feedback.
-Interestingly, we did not see large differences after removing chain-of-thought reasoning via test generation.
-In some cases, removing it even slightly improves the fault localization results.
-One reason may be that LLM4FL already includes self-reflection to make LLM think deeply about the result, and adding a chain-of-thought did not further improve the result.
 Our finding highlights the effectiveness of self-reflection, which should be considered in future fault localization results.
-The results show that each component of LLM4FL contributes to its overall fault localization performance, with coverage splitting and prompt chaining having the largest positive impact.
-Removing these components leads to significant declines in accuracy, confirming their crucial role in handling token limitations and enabling more effective multi-agent collaboration.
-4.3. RQ3: Does order matter in the initial list of methods provided to the LLM?
+The results show that each component of LLM4FL contributes to its overall fault localization performance, with coverage division and code navigation having the largest positive impact.
+Removing these components leads to significant declines in accuracy, confirming their role in finding faults in a large code-base.
+RQ3: Does order matter in the initial list of methods provided to the LLM?
 Motivation.
-LLM4FL divides the coverage data into segments to address the token size limitation of LLMs.
-We sort the methods using the Ochiai scores before segmentation, though different sorting mechanisms may affect the final fault localization result.
-Although LLM4FL eventually visits and assesses every method, a recent study (Chen et al., 2024) observes that the order of premises affects LLM’s results.
+LLM4FL divides the coverage data into different divisions using an order-aware division strategy to address the token size limitation of LLMs.
+We sort the methods using the Ochiai scores before the division, though different sorting mechanisms may affect the final fault localization result.
+Although LLM4FL eventually visits and assesses every method, a recent study (Chen et al., 2024a) observes that the order of premises affects LLM’s results.
 However, whether this effect extends to software engineering tasks, particularly fault localization, remains unclear.
-Hence, in this research question, we investigate whether the order of methods within the segments affects the LLM’s fault localization performance.
+Hence, in this research question, we investigate whether the order of methods within the groups affects the LLM’s fault localization performance.
 Approach.
-To test the effect of method ordering, we experiment with three distinct sorting strategies: LLM4FLExecution, LLM4FLOchiai (the default sorting in LLM4FL), and LLM4FLDepGraph.
-Each strategy provides a different way to sort the methods before they are segmented for further analysis.
-LLM4FLExecution: In this baseline approach, we use the unsorted list of methods executed during testing, as generated by Gzoltar (Campos et al., 2012).
+To test the effect of method ordering, we experiment with three distinct sorting strategies: LLM4FLExecution, LLM4FLOchiai (the default sorting in LLM4FL), and LLM4FLDepGraph to sort the methods before the divide-and-conquer step.
+LLM4FLExecution: We use the unsorted list of methods executed during testing, as generated by Gzoltar (Campos et al., 2012).
 This default list represents the natural execution order of the methods, with no explicit ranking or prioritization.
 By providing the LLM with methods based on the execution sequence, we establish a control case to measure its performance without any ranking influence.
-The result shows how the order of premises in reasoning tasks can lead to varied LLM performance in software engineering context (Chen et al., 2024).
-LLM4FLOchiai: As discussed in Section 3.1, we apply Ochiai to sort the methods during the segmentation process.
+LLM4FLOchiai: As discussed in Section 3.1, we apply Ochiai to sort the methods before the divide-and-conquer process.
 Ochiai is unsupervised and is efficient to compute.
 We hypothesize that providing the LLM with methods sorted by their suspiciousness score will lead to more effective fault localization, as the model will focus on the most likely faulty candidates earlier in the process.
-LLM4FLDepGraph: This approach uses the ranking produced by DepGraph, a state-of-the-art Graph Neural Network (GNN)-based fault localization technique (Li et al., 2015; Rafi et al., 2024), to sort the methods during the segmentation process.
+LLM4FLDepGraph: This approach uses the ranking produced by DepGraph, a state-of-the-art Graph Neural Network (GNN)-based fault localization technique (Li et al., 2015; Rafi et al., 2024), to sort the methods.
 DepGraph ranks methods based on structural code dependencies and code change history.
@@
 Figure 3 shows the fault localization results when using different sorting strategies.
-When methods were presented in an execution-based order, LLM4FLExecution achieved a Top-1 score of 299, gradually increasing to 402 for Top-3, 431 for Top-5, and reaching 462 for Top-10.
+When methods were presented in the execution order, LLM4FLExecution achieved a Top-1 score of 299, 402 for Top-3, 431 for Top-5, and 462 for Top-10.
 This performance establishes a baseline, showing how the LLM behaves without strategic ordering.
-However, sorting methods with the lightweight Ochiai scores resulted in noticeable improvements.
-LLM4FLOchiai improved the Top-1 score to 323, an 8% increase over LLM4FLExecution.
-The improvements also extended to other metrics, with the Top-3 score increasing to 421, the Top-5 reaching 469, and the Top-10 rising to 490.
+However, sorting methods with the lightweight Ochiai scores resulted in noticeable improvements across all Top-N, where LLM4FLOchiai improved the Top-1 score to 323, an 8% increase over LLM4FLExecution.
 LLM4FLDepGraph provides further improvement to the already-promising result of DepGraph, indicating method ordering is critical to LLM4FL, or LLM-based fault localization in general.
-LLM4FLDepGraph achieved the highest Top-1 score of 366, which is significantly outperforming both LLM4FLExecution and LLM4FLOchiai.
+LLM4FLDepGraph achieved the highest Top-1 score of 366, which significantly outperforms both LLM4FLExecution and LLM4FLOchiai.
 The improvement was consistent across all the metrics.
-For Top-3, the score rose to 492, providing a substantial boost over LLM4FLOchiai.
-Similarly, LLM4FLDepGraph excelled in the Top-5 and Top-10 categories, reaching 543 and 548, respectively.
 We also see that LLM4FLDepGraph has better Top-1, 3, and 5 scores compared to DepGraph.
 This consistent improvement underscores the importance of method ordering in enhancing the accuracy of LLM-based fault localization.
-This underscores the importance of method ordering for LLM-based fault localization, showing that lightweight techniques like Ochiai can achieve substantial performance gains without the heavy computational burden.
-Striking this balance between accuracy and efficiency makes these approaches particularly suitable for a wide range of fault localization tasks.
-However, it is important to recognize that LLM4FLDepGraph is computationally expensive due to its reliance on graph neural networks (GNNs).
-While LLM4FLDepGraph delivers top-tier results, LLM4FLOchiai also offers a strong, efficient alternative, delivering good localization accuracy at a fraction of the computational cost due to the unsupervised nature of Ochiai.
-In other words, LLM4FLOchiai may be more easily adapted in practice.
-Nevertheless, our finding establishes a new research direction for LLM-based fault localization.
-It demonstrates that intelligent method ordering strategies significantly impact the result of LLM-based fault localization.
-This approach opens up further opportunities for optimizing LLM-based fault localization by exploring more advanced ordering techniques and how different premises of ordering affect other software engineering tasks.
-The findings highlight that method ordering plays a crucial role in improving the performance of LLM-based fault localization, with a difference of up to 22% in Top-1 scores.
-While LLM4FLDepGraph delivers the best results, the lightweight LLM4FLOchiai offers a highly efficient alternative, providing significant accuracy gains with far lower computational costs, making it more practical for real-world adoption.
+Namely, if the initial order is closer to the group truth, the final localization result also tends to be more accurate.
+Our finding establishes a new research direction for LLM-based fault localization, or any software engineering tasks that take a list of software artifacts as input.
+Future studies may study how different premises of ordering affect other software engineering tasks, and how to combine traditional software engineering techniques to pre-process LLM’s input to further improve the results.
+The initial method ordering significantly impacts the accuracy of LLM-based fault localization, with Top-1 scores varying by up to 22%.
+Future research should explore different ordering strategies and how traditional software engineering techniques can be integrated to optimize LLM performance further.
 5. Threats to Validity
 Internal Validity.
-A potential threat to internal validity is the risk of data leakage in large language models (LLMs), where the model might have been exposed to data similar to the benchmark or specific fault localization cases during training.
-This could result in artificially inflated performance as the model may have prior knowledge of the evaluation data.
-Nevertheless, we mitigate this risk by preventing any content from being entered into ChatGPT that is related to the project name, the human-written bug report, or the bug ID.
+A potential threat to internal validity is the risk of data leakage in large language models (LLMs), where the model might have been exposed to the benchmark data during training.
+Nevertheless, Ramos et al.
+(Ramos et al., 2024) found that newer and larger models trained on larger datasets exhibit limited evidence of leakage for defect benchmarks.
+Since we utilize GPT-4o-mini, a large model trained using a tremendous amount of data, it likely shares similar characteristics in reducing memorization risks.
+Additionally, following prior work, we ensure no content related to the project name, human-written bug report, or bug ID is entered into ChatGPT to minimize the risk of data memorization (Qin et al., 2024).
 External Validity.
@@
 Our evaluation is based on Defects4J, a well-established dataset in the software engineering community.
-Although this dataset includes real-world bugs, the systems studied are primarily Java-based, and it remains uncertain whether our findings will generalize to other programming languages or domains.
+Although this dataset includes real-world bugs, the systems studied are primarily Java-based.
+Future studies may extend our study to other programming languages or domains.
 Construct Validity.
 Construct validity relates to whether the metrics we used accurately measure the performance of fault localization techniques.
-We used widely accepted Top-N metrics, which are commonly utilized in prior fault localization studies.
+We used widely accepted Top-N metrics commonly utilized in prior fault localization studies.
 However, our results are based on the assumption that developers primarily focus on the top-ranked faulty methods.
@@
 6. Conclusion
-In this paper, we introduced LLM4FL, an LLM-agent-based fault localization technique that addresses the challenges of traditional fault localization methods and the limitations of current LLM-based approaches.
-By combining SBFL rankings with a divide-and-conquer strategy, LLM4FL groups large coverage data into manageable pieces, enabling effective analysis within the token limitations of LLMs.
-The use of multiple agents, prompt chaining, and self-reflection techniques allows for iterative refinement of fault localization, while chain-of-thought reasoning further enhances accuracy by generating potential fixes and re-rank the methods.
-Our evaluation on the Defects4J (V2.0.0) benchmark demonstrated that LLM4FL outperforms existing LLM-based techniques, achieving 19.27% higher Top-1 accuracy compared to AutoFL, and surpasses supervised approaches such as DeepFL and Grace without task-specific training.
-We also found that key components like coverage splitting and prompt chaining are essential to LLM4FL’s success, with method ordering playing a significant role in performance, improving Top-1 accuracy by up to 22%.
-Overall, LLM4FL presents a scalable and efficient solution for fault localization, providing improved accuracy and reducing computational costs, making it practical for real-world software projects.
+In this paper, we introduced LLM4FL, an LLM-agent-based fault localization approach.
+It utilizes multiple specialized LLM agents, including Context, Debugger, and Reviewer, to iteratively refine and improve the accuracy of fault localization through order-aware prioritization, graph-based code navigation, and self-reflection using verbal reinforcement learning.
+Evaluated on the Defects4J (V2.0.0) benchmark, LLM4FL demonstrated significant improvements over existing approaches, achieving an 18.55% increase in Top-1 accuracy compared to AutoFL and 4.82% improvement over SoapFL.
+Further enhancements, including coverage segmentation and iterative refinement, increased accuracy by up to 22%.
 Future work will explore expanding LLM4FL’s capabilities for larger and more diverse codebases, further refining the agent collaboration and reasoning mechanisms.